### v1

# High-Throughput and Scalable Secure Inference Protocols for Deep Learning with Packed Secret Sharing

Abstract
Most existing secure neural network inference protocols based on secure multi-party computation (MPC) typically support at most four participants, demonstrating severely limited scalability. Liu et al. (USENIX Security’24) presented the first relatively practical approach by utilizing Shamir secret sharing with Mersenne prime fields. However, when processing deeper neural networks such as VGG16, their protocols incur substantial communication overhead, resulting in particularly significant latency in wide-area network (WAN) environments. In this paper, we propose a high-throughput and scalable MPC protocol for neural network inference against semi-honest adversaries in the honest-majority setting. The core of our approach lies in leveraging packed Shamir secret sharing (PSS) to enable parallel computation and reduce communication complexity. The main contributions are three-fold: i) We present a communication-efficient protocol for vector–matrix multiplication, based on our newly defined notion of vector–matrix multiplication-friendly random share tuples. ii) We design the filter packing approach that enables parallel convolution. iii) We further extend all non-linear protocols based on Shamir secret sharing to the PSS-based protocols for achieving parallel non-linear operations. Extensive experiments across various datasets and neural networks demonstrate the superiority of our approach in WAN. Compared to Liu et al. (USENIX Security’24), our scheme reduces the communication upto $5.85\times$, $11.17\times$, and $6.83\times$ in offline, online and total communication overhead, respectively. In addition, our scheme is upto $1.59\times$, $2.61\times$, and $1.75\times$ faster in offline, online and total running time, respectively.
I Introduction
Deep learning (DL) has emerged as a transformative methodology with widespread applications across diverse domains, particularly in computer vision and natural language processing. An increasing number of enterprises are deploying pre-trained DL models on computational servers to provide clients with inference services. However, this computational paradigm introduces significant privacy risks that warrant careful consideration. For example, the input data transmitted by clients may contain sensitive information, and DL models are often regarded as valuable assets that enterprises are reluctant to disclose. To address these dual privacy challenges, secure multi-party computation (MPC) is used to safeguard both input data and model privacy.
MPC enables a set of $n$ parties to mutually compute some function on their private inputs without compromising the privacy of these inputs or the correctness of the outputs [13, 14, 15, 16]. In recent years, numerous studies have been conducted to enhance the communication and computational efficiency of MPC-based secure neural network inference. In the dishonest-majority setting, most existing protocols focus on the two-party scenario [17, 20, 21, 6, 18, 22, 23, 24, 25, 26]. In the honest-majority setting, most existing protocols consider the three-party scenario [9, 12, 27, 28, 29, 30] and four-party scenario [31, 32, 33, 34], where the protocols tolerate one corrupted party. Unfortunately, all the aforementioned protocols universally suffer from scalability limitations. This critical shortcoming was recently addressed by Liu et al. [5], who designed a scalable protocol that can support a large number of parties for secure neural network inference based on Shamir secret sharing [2] against semi-honest adversaries in the honest-majority setting. Nevertheless, when we rerun the scalable protocol on relatively deeper neural networks such as VGG16, it exhibits poor performance owing to significant communication overhead. Hence, It is necessary to develop a more communication-efficient protocol that can better accommodate the requirements of real-world scenarios.
In theory, Franklin and Yung introduced a parallelizable technique called packed Shamir secret sharing (PSS) [1] where $k$ secrets are packed into a single secret sharing. Its follow-up works [8, 3, 48] extend functionality for secure computation. Particularly, a new technique called sharing transformation is proposed in work [3]. The technique allows us to perform arbitrary linear maps on the secrets of PSS with optimal communication complexity. Based on these techniques, we are able to achieve parallel secure inference.
Case I: For batch inference over $k$ input instances, naively upgrading the Shamir secret sharing scheme used in the protocol proposed by [5] to PSS could result in an amortized reduction in communication overhead by a factor of approximately $k$. However, in certain scenarios, such as a photo-based recognition application, only a single input may be available. If the above packed approach across multiple instances is still employed, it needs to introduce dummy inputs and pack them together with the real input to complete the inference process. More critically, this approach degenerates into a secure inference protocol similar to [5], thereby completely forfeiting the advantage of parallel computing provided by PSS.
Case II: Given that the method described in Case I isn’t efficient for processing the single input, we can turn to employ the substantial parallelism in a single inference. To be specific, deep neural networks typically consist of linear layers and non-linear layers. Linear layers primarily execute matrix operations. For example, the fully connected layer is a vector-matrix multiplication. The convolutional layer can be expressed as a matrix multiplication [12], which can be achieved through multiple vector-matrix multiplications. All these operations can be efficiently implemented using blocked matrix algorithms, which are particularly amenable to parallel computing acceleration. Under the block-processing strategy, each block of the vector and matrix is encoded into an individual PSS value, and block-wise multiply-accumulate operations are performed using DN protocol based on PSS (a natural extension of the DN protocol in [4]). Fig. 1 shows a toy example.
Regarding non-linear layers, such as the ReLU function defined as $\text{ReLU}(x)=\text{Max}(0,x)$, the operations are element-wise. This means each output value depends solely on the single value at the corresponding position in the input tensor. Hence, the parallel non-linear operations are entirely achievable by packing $k$ values into an individual PSS value. With input and output always kept as PSS, consecutive layers are composable; consequently, the complete secure-inference pipeline is executed in parallel.
Figure 1: A toy example. Boxes with the same color represent a packed secret sharing, in which $k$ values are packed.
I-A Practical Problem & Technique Overview
Nevertheless, a simple exploitation of parallelism in the inference process, as in Case II, is not practical, especially for linear layers.
First, the naive vector-matrix multiplication in Case II eliminates the benefit of using the PSS. As depicted in Fig. 1, each block of the vector is encoded into an individual PSS value, and block-wise multiply-accumulate operations are performed using DN protocol. Since the computed result remains in a partitioned form, this necessitates a sharing transformation to combine these partial results into the final output. Compared with Shamir secret sharing-based approach [5], the entire process instead incurs more communication overhead. The reason is that [5] only involves the DN protocol for vector-matrix multiplication.
Second, when processing convolutional layers, zero padding of convolutional inputs incurs significant communication cost. As depicted in Case II, a convolutional layer can be achieved through multiple vector-matrix multiplications. In addition, zero padding is aslo frequently applied to the convolutional inputs in many neural networks, such as VGG16. Intuitively, we can directly achieve convolution computation by the naive vector-matrix multiplication. Unfortunately, when zero padding needs to be applied to convolutional inputs, it becomes necessary to combine all inputs with these zero padding values into new PSS values. This combination process incurs significant communication overhead.
To address the first issue, inspired by the sharing transformation technique, we first define vector-matrix multiplication-friendly random share tuples (VM-RandTuple), which can be generated in batches by employing the Vandermonde matrix in the offline phase. Furthermore, in the DN protocol, the parties locally compute and send the result to $P_{1}$. Then, $P_{1}$ reconstructs the result and sends it back to the parties. A simple optimization to the second step is having $P_{1}$ combine these partial results into the final output and reshare the final output to the parties. Theoretically, integrating VM-RandTuple and the optimized DN protocol has the potential to achieve better online communication complexity vector-matrix multiplication.
To address the second issue, we observed that the convolutional input is required to interact with every filter. Hence, the convolution process inherently involves parallel computation. In light of this, we propose a filter packing approach, i.e., we pack $k$ values that are selected respectively from the same position of $k$ filters into a single packed secret share. Correspondingly, we also pack $k$ copies of the input into a single packed secret share. Then based on the packing approach, we achieve completing the convolution computation in parallel. Further, when zero padding is applied to the convolutional input without communication. Finally, since $k$ copies of the input are packed into a single packed secret share, we also carefully account for the inter-layer connectivity by designing some protocols.
I-B Contributions
Building upon the above techniques, in this paper, we propose a parallel and scalable MPC protocol by leveraging PSS for secure inference, where the model from companies and input data from clients are secret-shared among all computational servers. The main contributions are summarized below.
  • Efficient Vector-matrix Multiplication. We define vector-matrix multiplication-friendly random share tuples and design a protocol to generate these tuples for computing vector-matrix multiplication. We further combine these tuples with the DN [4] protocol, resulting in an efficient protocol for vector-matrix multiplication with a single round of online communication.
  • Parallel Convolution Computation. We propose a filter packing approach for convolution computation. By integrating our method with the DN protocol, we achieve a parallel convolution computation protocol that requires merely single-round online communication. Our convolution computation protocol can further support zero padding without extra communication.
  • Parallel Non-linear Operations. We extend all protocols based on Shamir secret sharing for non-linear operations in [5] to the PSS-based protocols, which enable parallel computation of the non-linear operations, resulting in efficient protocols for parallel ReLU with $O(\text{log}_{2}{\ell})$ rounds of online communication.
  • Efficient Secure Inference. We implement all protocols and conduct extensive secure inference experiments on various neural networks in both LAN and WAN settings with varying numbers of parties and different packaging parameters. Experimental results demonstrate progressively enhanced performance advantages with increasing numbers of participating entities and deeper neural network architectures. Compared to Liu et al. [5], our scheme reduces the communication upto $5.85\times$, $11.17\times$, and $6.83\times$ in offline, online, and total communication overhead, respectively. At the same time, our scheme is upto $1.59\times$, $2.61\times$, and $1.75\times$ in offline, online and total running time, respectively. Furthermore, experimental results reveal the superiority of our protocol in terms of memory overhead. For example, we simulated 63 participants on 11 computational servers, each of which was equipped with 64GB of RAM. Our protocol reliably supports VGG16 inference. In contrast,  [5] suffers from memory overflow problems.
II Preliminary
II-A Notation
Let $p$ be a Mersenne prime, which is the form of $p=2^{\ell}-1$ for some prime $\ell$. In secure neural network inference, $\ell$ is chosen to be 31 and 61, which is sufficient. $\mathbb{F}_{p}$ denotes a finite field. We denote by $[n]$ the set $\{0,1,\cdots,n-1\}$, where $n\in\mathbb{N},\mathbb{N}$ is the set of natural numbers. We use the lowercase letter $x$ to denote a scalar and $x_{i}$ to denote the $i$-th bit of the scalar $x$. The bold lowercase letter $\boldsymbol{x}=\{\boldsymbol{x_{i}}\}_{i=0}^{n-1}$ is a vector, $\boldsymbol{x}_{i}$ is the $i$-th element of the vector $\boldsymbol{x}$, $\boldsymbol{x}_{i,j}$ is $j$-th bit of the $i$-th element of the vector $\boldsymbol{x}$ and $\boldsymbol{x}^{i}$ is indexed by $i$ to distinguish different vectors. Correspondingly, uppercase letter $X$ is a matrix and $X_{i,j}$ is the element indexed by row $i$ and column $j$ of the matrix $X$. The bold uppercase letter $\boldsymbol{X}$ is a tensor. Van(n,nt)Van(n,n-t) is an $n$-by-(nt)(n-t) Vandermonde matrix.
II-B Packed Shamir Secret Sharing Scheme
The packed Shamir secret sharing (PSS) scheme [1] is a natural generalization of the standard Shamir secret sharing scheme [2].
For a vector $\boldsymbol{x}\in\mathbb{F}_{q}^{k}$, we will use $\llbracket\boldsymbol{x}\rrbracket_{d}\in\mathbb{F}_{p}$ to denote a degree-$d$ PSS, where $k-1\leq d\leq n-1$. It requires $d+1$ shares to reconstruct the whole secrets, and any dk+1d-k+1 shares are independent of the secrets. The scheme $\llbracket\boldsymbol{x}\rrbracket_{d}$ corresponds to a degree-$d$ polynomial $f$ such that the values $\{f(i)\}_{i=1}^{n}$ represent the shares and $\{f(\boldsymbol{s}_{i})=\boldsymbol{x}_{i}\}_{i=0}^{k-1}$, where $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ are the default positions to store the secrets $\boldsymbol{x}=\{\boldsymbol{x}_{i}\}_{i=0}^{k-1}$. The PSS scheme has the following nice properties:
  • Linear Homomorphism: For all $d\geq k-1$ and $\boldsymbol{x},\boldsymbol{y}\in\mathbb{F}_{q}^{k},\llbracket\boldsymbol{x}\rrbracket_{d}+\llbracket\boldsymbol{y}\rrbracket_{d}=\llbracket\boldsymbol{x}+\boldsymbol{y}\rrbracket_{d}$. Assume that $c$ is a public constant, $c+\llbracket\boldsymbol{x}\rrbracket_{d}$ and $c\times\llbracket\boldsymbol{x}\rrbracket_{d}$ denote adding $c$ to each element of $\boldsymbol{x}$ and multiplying each element of $\boldsymbol{x}$ by $c$, respectively.
  • Multiplication-friendliness in [3]: When $d\leq n-k$, all parties can locally multiply a public vector $\boldsymbol{c}\in\mathbb{F}_{p}^{k}$ with the share $\llbracket\boldsymbol{x}\rrbracket_{d}$. First, all parties locally compute $\llbracket\boldsymbol{c}\rrbracket_{k-1}$. Then all parties can locally compute $\llbracket\boldsymbol{c}*\boldsymbol{x}\rrbracket_{k-1+d}=\llbracket\boldsymbol{c}\rrbracket_{k-1}\times\llbracket\boldsymbol{x}\rrbracket_{d}$, where * denote coordinate-wise multiplication.
  • Multiplicative: For all $d_{1},d_{2}\geq k-1,d_{1}+d_{2}<n$ and for all $\boldsymbol{x},\boldsymbol{y}\in\mathbb{F}_{q}^{k},\llbracket\boldsymbol{x}\rrbracket_{d_{1}}\times\llbracket\boldsymbol{y}\rrbracket_{d_{2}}=\llbracket\boldsymbol{x}*\boldsymbol{y}\rrbracket_{d_{1}+d_{2}}$. If $d_{1}=d_{2}=\frac{n-1}{2}$, we can compute $\llbracket\boldsymbol{x}\rrbracket_{d}*\llbracket\boldsymbol{y}\rrbracket_{d}=\llbracket\boldsymbol{x}\times\boldsymbol{y}\rrbracket_{d}$ by a natural extension of the DN multiplication protocol in [4]. We summarize the functionality $\mathcal{F}_{Pmult-DN}$ in Appendix -A. The protocol is as follows:
    • (1) There are $n$ parties $\{P_{i}\}_{i=1}^{n}$. They prepare a pair of two random sharings ($\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r}\rrbracket_{d}$) of the same secrets vector $\boldsymbol{r}\in\mathbb{F}_{p}^{k}$.
    • (2) All parties compute $\llbracket\boldsymbol{a}\rrbracket_{2d}=\llbracket\boldsymbol{x}\rrbracket_{d}\times\llbracket\boldsymbol{y}\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{2d}$ and send $\llbracket\boldsymbol{a}\rrbracket_{2d}$ to $P_{1}$.
    • (3) $P_{1}$ reconstructs the secrets $\boldsymbol{a}$, generates $\llbracket\boldsymbol{a}\rrbracket_{d}$, and distributes the shares to all other parties.
    • (4) All parties compute $\llbracket\boldsymbol{z}\rrbracket_{d}=\llbracket\boldsymbol{a}\rrbracket_{d}-\llbracket\boldsymbol{r}\rrbracket_{d}$.
In addition, a degree-$t$ Shamir secret sharing scheme corresponds to a degree-$t$ polynomial $f$ such that $\{f(i)\}_{i=1}^{n}$ are the shares and $f(0)$ is the secret. However, we do not always need to store the secret at point 0. We use $x|_{j}]_{t}$ to denote a degree-$t$ Shamir sharing whose secret is stored at the point $j,j\notin\{1,\cdots,n\}$.
Now let $E_{i}\in\{0,1\}^{k}$ denote the unit vector, whose $i$-th entry is 1 and all other entries are 0. Then parties can convert $E_{i}$ to $\llbracket E_{i}\rrbracket_{k-1}$ locally. There are the following nice properties:
  • Transforming Shamir secret shares at $k$ distinct positions into a single PSS locally:
    $\llbracket\boldsymbol{x}\rrbracket_{t+k-1}=\sum_{i=0}^{k-1}\llbracket E_{i}\rrbracket_{k-1}\cdot[\boldsymbol{x}_{i}|_{s_{i}}]_{t}$.
  • Correspondingly, we can choose $k$ secrets from different position of $k$ distinct PSS, and generates a new PSS that the positions of the secrets we choose are from different positions remain unchanged in the new PSS:
    $\llbracket\boldsymbol{x}\rrbracket_{d+k-1}=\sum_{i=0}^{k-1}\llbracket E_{i}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{x}^{i}\rrbracket_{d}$,
    where $\boldsymbol{x}=\{\boldsymbol{x}^{i}_{i}\}_{i=0}^{k-1}$.
II-C Neural Network
In this paper, we focus on the convolutional neural networks (CNN), a widely used in computer vision. It mainly consists of fully connected layers, convolutional layers and non-linear layers.
Fully Connected Layer (FC). The input and output of FC are vector $\boldsymbol{v}\in\mathbb{F}_{p}^{n_{i}}$ and $\boldsymbol{u}\in\mathbb{F}_{p}^{n_{0}}$ respectively. An FC is formalized as a tuple $(W,\boldsymbol{b})$ where $W$ is a $(n_{i}\times n_{o})$ weight matrix and $\boldsymbol{b}$ is a bias vector of length $n_{o}$. The output is represented as a linear transformation $\boldsymbol{u}=\boldsymbol{v}W+\boldsymbol{b}$. In a word, the FC is a vector-matrix multiplication. We propose an efficient protocol for computing vector-matrix multiplication in §III.
Convolution Layer (Conv). The input to a Conv layer is a tensor $\boldsymbol{X}\in\mathbb{R}^{w_{i}\times h_{i}\times c_{i}}$ where $w$ is the image width, $h$ is the image height, and $c$ is the number of input channels. The Conv is then parameterized by the tuple $(f_{w},f_{h},c_{i},c_{o})$. In other words, there are $c_{o}$ filter banks, each consisting of $c_{i}$ many $f_{w}\times f_{h}$ filters. The Conv process is sliding each filter with a certain stride $s$ on the tensor $\boldsymbol{X}$. The output of the Conv can then be parameterized by the tuple $(w_{o},h_{o},c_{o})$ where $w_{o}=\lfloor(w_{i}-f_{w}+2p)/s\rfloor+1,h_{o}=\lfloor(h_{i}-f_{h}+2p)/s\rfloor+1,p$ is padding size applied to the edges of each $w_{i}\times h_{i}$ matrix of the tensor $\boldsymbol{X}$. Fig. 2 shows the Conv process. For efficiently achieving convolution computation, we design an efficient protocol in §IV.
Figure 2: Convolution process.
Non-Linear Layer (ReLU and Pool). In deep learning, the non-linear layers consist of activation functions and pooling functions. The activation functions are applied in each element of input independently. The common activation functions is ReLU function: $\text{ReLU}(x)=\text{Max}(0,x)$. The pooling function reduces the output size. One of the most popular pooling functions in CNN is the MaxPool function that arranges inputs into several windows and calculates the maximum for each window. We construct the parallel non-linear operators in §V.
II-D Model
We use an outsourcing computing model for secure inference. We assume there are a client, a model owner and $n\in\mathbb{N}$ computational servers, denoted by $P_{1},P_{2},\cdots,P_{n},$ where $n=2d+1,d\in\mathbb{N}$. When a client wants to perform secure inference, the client and model owner respectively distribute their own data across $n$ servers by a degree-$d$ PSS scheme. Then $n$ servers complete a secure inference and send the result shares to the client. Finally, the client can get the inference result by reconstructing the shares.
We describe our protocols in the semi-honest model with passive security. The client and model owner are honest. The number of corrupted servers that are assumed to be poly-time bounded is at most t=dk+1t=d-k+1 for a natural extension of the DN [4] multiplication protocol, where $2\leq k\leq d$. If $k=1$, the PSS becomes standard Shamir secret sharing. The client and the model owner do not collude with the servers. It is worth noting that a degree-$d$ PSS can pack $k\in\mathbb{N}$ secrets, and $k$ and default positions $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ are publicly known to the client, model owner and servers.
III Efficient Vector-Matrix Multiplication
In this section, we first present the design of the vector-matrix multiplication protocol in §III-A. Then, we extend the vector-matrix multiplication protocol to support truncation in §III-B.
Figure 3: A toy example. Boxes with the same color represent a packed secret sharing, in which $k$ values are packed and located at positions $(\boldsymbol{s}_{0},\cdots,\boldsymbol{s}_{k-1})$; blank boxes indicate plaintext values. ① and ② correspond to general vector-matrix multiplication. ① and ③ correspond to fixed-point vector-matrix multiplication.
III-A Protocol Design
FC is a key layer in the deep learning §II-C, which consists of vector-matrix multiplications. In light of the earlier analysis, vector-matrix multiplication $\boldsymbol{a}\cdot A$ can be calculated by blocked matrix algorithms, where $\boldsymbol{a}\in\mathbb{F}_{p}^{u},A\in\mathbb{F}_{p}^{u\times v}$. Specifically, we divide the vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ into $\lceil u/k\rceil$ blocks and divide each column of the matrix $A\in\mathbb{F}_{p}^{u\times v}$ into $\lceil u/k\rceil$ blocks. The size of each block is $k$. Then we pack each block into a PSS (e.g., $\llbracket\boldsymbol{a}\rrbracket\in\mathbb{F}_{p}^{\lceil u/k\rceil},\llbracket A\rrbracket\in\mathbb{F}_{p}^{\lceil u/k\rceil\times v}$) and combine it with the DN [4] protocol.
In order to reduce communication cost, we first define vector-matrix multiplication-friendly random share tuple (VM-RandTuple), i.e., $(\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{k},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p})$, where $\boldsymbol{r}=\{\boldsymbol{r}_{i}\}_{i=0}^{k^{2}-1},\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}\}_{i=0}^{k-1}$. We can utilize a Vandermonde matrix for the batch generation of such random shares tuple in the offline phase. We formally define the functionality $\mathcal{F}_{VM-RandTuple}$ in Appendix -A and provide a secure protocol implementation in Algorithm 1. Based on the VM-RandTuple, a key optimization to the second step of DN protocol is having $P_{1}$ combine these partial results into the final output and reshare the final output to the parties. This optimization directly leads to a $k$-fold reduction in the amortized communication cost of the second step of DN protocol, compared with [5].
We describe a toy example in Fig. 3. It embodies a core idea of our technology. ① and ② correspond to general vector-matrix multiplication; ① and ③ correspond to fixed-point vector-matrix multiplication. We defer the detailed the latter to §III-B. In Fig. 3, we set $k=2$, meaning that $2$ values are packed into a single packed secret share. Accordingly, the vector $\boldsymbol{a}$ is packed into 2 packed secret shares, i.e., $\llbracket\boldsymbol{a}\rrbracket\in\mathbb{F}_{p}^{2}$ and the matrix $A$ is packed into $2\times 2$ packed secret shares, i.e., $\llbracket A\rrbracket\in\mathbb{F}_{p}^{2\times 2}$. Then all parties $\{P_{i}\}_{i=1}^{n}$ can compute $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}=\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}$ locally. Since $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}$ has a degree of $2d$, a degree reduction is required. $P_{i}$ sends the $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}+\llbracket\boldsymbol{r}\rrbracket_{2d}$ to $P_{1}$ and $P_{1}$ reconstructs the secret values at the default positions $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ and then calculates the sum of the results. Finally, $P_{1}$ packs the multiple results, i.e., $(\boldsymbol{c}_{1},\boldsymbol{c}_{2})$ into a single degree-$d$ PSS and shares to the other parties. We formally define the functionality $\mathcal{F}_{VecMatMult}$ in Appendix -A and provide a secure protocol implementation in Algorithm 2.
To sum up, given a vector of length $u$ and a $u$-by-$v$ matrix, the online complexity is exactly 1 round with $(1+\frac{1}{k})uv$ field elements per party. The offline complexity is exactly 2 round with $(1+\frac{1}{k})\frac{2nv}{n+2k-1}$ field elements per party.
Input: None.
Output: $\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{(n-t)\cdot k},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p}^{(n-t)}$, where
$\boldsymbol{r}=\{\boldsymbol{r}_{i}\}_{i=0}^{(n-t)\cdot k^{2}-1},$ $\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}\}_{i=0}^{(n-t)\cdot k-1}.$
  1. 1.
    $P_{i},i\in\{1,\cdots,n\}$ randomly samples a vector $\boldsymbol{a}^{i}$ with length of $k^{2}$ and distributes $\llbracket\boldsymbol{a}^{i}\rrbracket_{2d}\in\mathbb{F}_{p}^{k}.$
  2. 2.
    $P_{i}$ computes $\boldsymbol{b}^{i}=\{\boldsymbol{b}_{u}^{i}=\sum_{j=u\cdot k}^{(u+1)\cdot k-1}\boldsymbol{a}^{i}_{j}\}_{u=0}^{k-1}$ and distributes $\boldsymbol{c}^{i}=\{[\boldsymbol{b}^{i}_{u}|_{\boldsymbol{s}_{u}}]_{t}\}_{u=0}^{k-1}.$
  3. 3.
    $P_{i}$ computes
    $A=Van(n,n-t)^{T}\cdot(\llbracket\boldsymbol{a}^{1}\rrbracket_{2d},\cdots,\llbracket\boldsymbol{a}^{n}\rrbracket_{2d}).$
    $B=Van(n,n-t)^{T}\cdot(\boldsymbol{c}^{1},\cdots,\boldsymbol{c}^{n}).$
  4. 4.
    for $j\leftarrow 0$ to nt1n-t-1 do
   Compute $\boldsymbol{f}_{j}=\sum_{i=0}^{k-1}\llbracket E_{i}\rrbracket_{k-1}\cdot B_{j,i}.$ for $u\leftarrow 0$ to k1k-1 do
      Set $\boldsymbol{g}_{j\cdot k+u}=A_{j,u}.$
  • 5.
    $P_{i}$ sets $\llbracket\boldsymbol{r}\rrbracket_{2d}=\boldsymbol{g}=\{\boldsymbol{g}_{i}\}_{i=0}^{(n-t)\cdot k},\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}=\boldsymbol{f}=\{\boldsymbol{f}_{i}\}_{i=0}^{n-t-1}.$ return $\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}.$
  • Algorithm 1 VM-RandTuple $\Pi_{VM-RandTuple}$
    There are $n$ parties where $n=2d+1,1<k\leq d$ and the number of corrupted servers is t=dk+1t=d-k+1. The original vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ and matrix $A\in\mathbb{F}_{p}^{u\times v}$.
    Input: $\llbracket\boldsymbol{a}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil u/k\rceil},\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil u/k\rceil\times v}$.
    Output: $\llbracket\boldsymbol{c}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil v/k\rceil},$ where $\boldsymbol{c}=\boldsymbol{a}\cdot A$.
    if $u,v$ is not divisible by $k$, the original vector and matrix can be zero-padded accordingly.
    Common Randomness: $(\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{v},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil v/k\rceil})\leftarrow\mathcal{F}_{VM-RandTuple}$ where $\boldsymbol{r}=\{\boldsymbol{r}_{i}\}^{kv-1}_{i=0},\boldsymbol{r^{\prime}}=\{\boldsymbol{r}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}\}^{v-1}_{i=0}.$
    1. 1.
      The parties compute $\llbracket\boldsymbol{z}\rrbracket_{2d}=\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{2d}$ and send $\llbracket\boldsymbol{z}\rrbracket_{2d}$ to the $P_{1}$.
    2. 2.
      $P_{1}$ reconstructs the values $\{\boldsymbol{z}_{i}\}_{i=0}^{kv-1}$, computes $\boldsymbol{f}=\{\boldsymbol{f}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{z}_{j}\}^{v-1}_{i=0}$, and distributs $\llbracket\boldsymbol{f}\rrbracket_{d}\in\mathbb{F}_{d}^{\lceil v/k\rceil}$ to other parties.
    3. 3.
      The parties compute $\llbracket\boldsymbol{c}\rrbracket_{d}=\llbracket\boldsymbol{f}\rrbracket_{d}-\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}.$
    return $\llbracket\boldsymbol{c}\rrbracket_{d}.$
    Algorithm 2 Vector-Matrix Mult $\Pi_{VecMatMult}$
    Theorem 1.
    $\Pi_{VM-RandTuple}$ securely realizes $\mathcal{F}_{VM-RandTuple}$ in the stand-alone model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Theorem 2.
    $\Pi_{VecMatMult}$ securely realizes $\mathcal{F}_{VecMatMult}$ in the $\mathcal{F}_{VM-RandTuple}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    III-B Truncation
    In secure inference, decimal numbers are encoded using fixed-point representation universally. In this representation, a decimal value is represented as $\ell$-bit integers, where the first $\ell-\ell_{x}$ bits are the integer part and the last $\ell_{x}$ bits are the fractional part. Multiplying two fixed-point numbers yields $2\ell_{x}$ fractional bits and $\ell-2\ell_{x}$ integer bits; truncation is required to reduce the fractional part back to $\ell_{x}$ bits, ensuring precision and avoiding overflow. [5] proposes a novel truncation protocol in the Mersenne prime field that reduces the gap between shares and the secret to just 1 bit. However, the new protocol is not well-suited for our work, as it incurs significant communication overhead. Specifically, the online communication overhead of the truncation protocol is the same as DN multiplication, but applying it in our work requires additional open operations and degree reduction. Therefore, we propose our fixed-point vector-matrix multiplication protocol that is a natural extension of ABY3 [9] in Algorithm 3. In Fig. 3, ① and ③ correspond to fixed-point vector-matrix multiplication.
    The original vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ and matrix $A\in\mathbb{F}_{p}^{u\times v}$.
    Input: $\llbracket\boldsymbol{a}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil u/k\rceil},\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil u/k\rceil\times v}$.
    Output: $\llbracket\boldsymbol{c}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil v/k\rceil},$ where $\boldsymbol{c}=\boldsymbol{a}\cdot A$.
    if $u,v$ is not divisible by $k$, the original vector and matrix can be zero-padded accordingly.
    Common Randomness: $(\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{v},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil v/k\rceil})\leftarrow\mathcal{F}_{TruncTriple}$ where $\boldsymbol{r}=\{\boldsymbol{r}_{i}\}^{kv-1}_{i=0},\boldsymbol{r^{\prime}}=\{\boldsymbol{r}^{\prime}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}/2^{\ell_{x}}\}^{v-1}_{i=0}.$
    1. 1.
      The parties compute $\llbracket\boldsymbol{z}\rrbracket_{2d}=\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{2d}$ and send $\llbracket\boldsymbol{z}\rrbracket_{2d}$ to the $P_{1}$.
    2. 2.
      $P_{1}$ reconstructs the values $\{\boldsymbol{z}_{i}\}_{i=0}^{kv-1}$, computes $\boldsymbol{f}=\{\boldsymbol{f}_{i}=(\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{z}_{j})/2^{\ell_{x}}\}^{v-1}_{i=0}$ and distributs $\llbracket\boldsymbol{f}\rrbracket_{d}\in\mathbb{F}_{d}^{\lceil v/k\rceil}$ to other parties.
    3. 3.
      The parties compute $\llbracket\boldsymbol{c}\rrbracket_{d}=\llbracket\boldsymbol{f}\rrbracket_{d}-\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}.$
    return $\llbracket\boldsymbol{c}\rrbracket_{d}.$
    Algorithm 3 Vector-Matrix Mult with Trunc $\Pi_{VecMatMult}^{Trunc}$
    We can see clearly that the key is how to generate the random shares $\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$. The vectors $\boldsymbol{r}^{\prime}$ and $\boldsymbol{r}$ satisfy the following relations: $\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}\}^{k-1}_{i=0},\boldsymbol{r}=\{\boldsymbol{r}_{i}\}^{k^{2}-1}_{i=0},\boldsymbol{r}^{\prime}_{i}=(\sum_{i=j\cdot k}^{(j+1)\cdot k-1}\boldsymbol{r}_{i})/2^{\ell_{x}},j\in[k]$. We formally define the functionality $\mathcal{F}_{TruncTriple}$ in Appendix -A and provide a secure protocol implementation in Algorithm 4. Specifically, $\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$ and $k$ random packed secret shares $\llbracket\boldsymbol{w}^{i}\rrbracket_{d},i\in[k]$ are genereted, where the length of vector $\boldsymbol{w}^{i}$ is $k$. Then there is the following relation: $\boldsymbol{r}_{j}=\boldsymbol{w}_{\lfloor j/k\rfloor}^{j\pmod{k}},j\in[k^{2}]$. Fig. 4 depicts a sample example. It should be noted that $\llbracket\boldsymbol{w}^{i}\rrbracket_{d}$ needs to be recombined. The recombination process consists of two steps: i) All secret values in vector $\boldsymbol{w}^{i}$ are converted into Shamir secret shares $\boldsymbol{w}^{i}_{j}|_{i}]_{d}$. Inspired by [10], we obtain Theorem 4. We treat a single packed secret share as $k$ distinct Shamir secret shares and apply Theorem 4 to achieve local conversion. We formalize this process in Algorithm 5; ii) These transformed values are recombined through linear combination with vector $\llbracket E_{i}\rrbracket_{k-1}$. Although Algorithm 5 performs the conversion of secret shares locally, it results in an increased degree of $2d$. Therefore, a degree reduction is required before recombination by functionality $\mathcal{F}_{DgreeTrans}$. We provide secure protocol implementation for $\mathcal{F}_{RandomBits},\mathcal{F}_{Random},\mathcal{F}_{DegreeTrans}$ in Appendix -B.
    Theorem 3.
    $\Pi_{VecMatMult}^{Trunc}$ securely realizes $\mathcal{F}_{VecMatMult}^{Trunc}$ in the $\mathcal{F}_{TruncTriple}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Theorem 4.
    A Shamir secret share $x|_{a}]_{d},$ where $f(a)=x,n=2d+1,a\notin\{1,2,\cdots,n\},f$ is a degree-$d$ polynomial. Each of the $n$ parties holds its own share $f(i),i\in\{1,2,\cdots,n\}$. The secret value $x$ can be locally converted to $x|_{b}]_{2d}$, with the conversion performed as $f^{\prime}(i)=f(i)\Pi_{j=1,j\neq i}^{n}\frac{a-j}{b-j}$.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Input: None.
    Output: $\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{k},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p}$,
    where $\boldsymbol{r}=\{\boldsymbol{r}_{i}\}^{k^{2}-1}_{i=0},$
    $\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}/2^{\ell_{x}}\}^{k-1}_{i=0}.$
    1. 1.
      The parties generete $\ell$ random bits PSS $\llbracket\boldsymbol{b}^{i}\rrbracket_{d}\leftarrow\mathcal{F}_{RandomBits}$, where the length of each vector $\boldsymbol{b}^{i}$ is $k,i\in[\ell]$.
    2. 2.
      Compute $\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}=\sum_{i=\ell_{x}}^{\ell-1}2^{i-\ell_{x}}\cdot\llbracket\boldsymbol{b}^{i}\rrbracket_{d},\llbracket\boldsymbol{q}\rrbracket_{d}=\sum_{i=0}^{\ell-1}2^{i}\cdot\llbracket\boldsymbol{b}^{i}\rrbracket_{d}.$
    3. 3.
      Generete k1k-1 random PSS $\llbracket\boldsymbol{w}^{i}\rrbracket_{d}\leftarrow\mathcal{F}_{Random},i\in[1,k-1]$, and compute $\llbracket\boldsymbol{w}^{0}\rrbracket_{d}=\llbracket\boldsymbol{q}\rrbracket_{d}-\sum_{i=1}^{k-1}\llbracket\boldsymbol{w}^{i}\rrbracket_{d}$, where the bit length of each element $\boldsymbol{w}^{i}_{j}$ is $\ell$.
    4. 4.
      Get $\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d},i,j\in[k]$ by ShConvert($\llbracket\boldsymbol{w}^{i}\rrbracket_{d}$) (Algorithm 5) locally and get $\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d-k+1}\leftarrow\mathcal{F}_{DegreeTrans}([\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d})$.
    5. 5.
      Compute $\llbracket\boldsymbol{r}\rrbracket_{2d}=\sum_{i=0}^{k-1}\llbracket E_{i}\rrbracket_{k-1}\cdot[\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d-k+1}.$
    return $\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}.$
    Algorithm 4 Trunc Triple $\Pi_{TruncTriple}$
    Theorem 5.
    $\Pi_{TruncTriple}$ securely realizes $\mathcal{F}_{TruncTriple}$ in the $(\mathcal{F}_{RandomBits},\mathcal{F}_{Random},\mathcal{F}_{DegreeTrans})$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Input: $\llbracket\boldsymbol{x}\rrbracket_{d},$ where $\llbracket\boldsymbol{x}\rrbracket_{d}$ can be regarded as $k$ distinct Shamir secret shares at position $\boldsymbol{s}_{0}$ to $\boldsymbol{s}_{k-1}$.
    Output: $\boldsymbol{x}_{v}|_{\boldsymbol{s}_{v}}]_{2d},v\in[k]$.
    1. 1.
      Each party computes locally for $v\leftarrow 0$ to k1k-1 do
     $\boldsymbol{x}_{v}|_{b}]_{2d}\leftarrow[\boldsymbol{x}_{v}|_{\boldsymbol{s}_{v}}]_{d}\cdot\Pi_{j=1,j\neq i}^{n}\frac{\boldsymbol{s}_{v}-j}{b-j},i\in[1,n]$;
    return $\boldsymbol{x}_{v}|_{b}]_{2d}.$
    Algorithm 5 ShConvert
    Figure 4: A sample example. $\boldsymbol{q}_{i}/2^{\ell_{x}}=\boldsymbol{r}^{\prime}_{i},i\in[3],\boldsymbol{q}_{0}=(\boldsymbol{r}_{0}+\boldsymbol{r}_{1}+\boldsymbol{r}_{2}),\boldsymbol{q}_{1}=(\boldsymbol{r}_{3}+\boldsymbol{r}_{4}+\boldsymbol{r}_{5}),\boldsymbol{q}_{2}=(\boldsymbol{r}_{6}+\boldsymbol{r}_{7}+\boldsymbol{r}_{8}).$
    To sum up, given a vector of length $u$ and a $u$-by-$v$ matrix, the online complexity is the same as protocol $\Pi_{VecMatMult}$. The offline complexity of this protocol depends on the generation of the random bits, random values and degree transformation, that is 5 rounds and $(n-1)(\frac{4\ell+k\ell+2k^{2}}{n-t}+\frac{4\ell+2k^{2}}{n}+k\ell+\ell)$ field elements per party for a truncation triple.
    IV Efficient Convolutions
    In this section, we first present the design of the convolution computation protocol in §IV-A. Then, we introduce the connections between layers in §IV-B.
    IV-A Protocol Design
    FC is exactly a vector-matrix multiplication. Similarly, a Conv §II-C can also be expressed as a matrix multiplication that can be achieved through multiple vector-matrix multiplications. In light of the above analysis, when zero padding needs to be applied to convolutional input, the combination process incurs significant communication overhead.
    For avoiding significant communication cost, we present our efficient convolution computation protocol by leveraging the inherent parallelizability of convolution computation. To be specific, the input tensor or image is required to interact with every filter as shown in Fig. 2. Hence, the convolution process inherently involves parallel computation. In light of this, we propose a filter packing approach, i.e., we pack $k$ values that are selected respectively from the same position of $k$ filters into a single packed secret share. Correspondingly, we also pack $k$ copies of the input into a single packed secret share. Fig. 5 depicts a sample example. Then we can complete convolution computation by using a parallel matrix multiplication protocol. Further, when zero padding is applied to the input of convolutional operations, it can be performed locally without communication since we can directly pack $k$ zero values into a single packed secret share.
    We directly present the protocol for parallel fixed-point matrix multiplication in Algorithm 6 and formally define the functionality $\mathcal{F}_{PMatMult}^{Trunc}$ in Appendix -A. In offline Step 2, it holds $R_{\alpha,\beta}=\sum_{i=0}^{\ell-1}2^{i}\cdot R^{\prime\prime i}_{\alpha,\beta}\cdot R^{\prime\prime i}_{\alpha,\beta},$ since $R^{\prime\prime i}_{\alpha,\beta}\in\{0,1\}^{k}.$ In addition, $\llbracket\boldsymbol{0}\rrbracket_{d}$ can be generated by PRG without communication.
    To sum up, given a $u$-by-$v$ matrix and a $v$-by-$m$ matrix, the online complexity is 1 round with $\frac{2um}{k}$ field elements per party. The offline complexity is 4 rounds and $\frac{um\ell}{k}(\frac{3n}{n+2k-1}+2+\frac{n+2k}{2n})$ field elements per party.
    $2k$ original matrices $A^{i}\in\mathbb{F}_{p}^{u\times v},B^{i}\in\mathbb{F}_{p}^{v\times m},i\in[k]$. $\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{u\times v}$ and $\llbracket B\rrbracket_{d}\in\mathbb{F}_{p}^{v\times m}$ pack all $A^{i}$ and all $B^{i}$, respectively. All $k$ $A^{i}_{\alpha,\beta},\alpha\in[u],\beta\in[v]$ are packed into a packed secret share $\llbracket A_{\alpha,\beta}\rrbracket_{d}\in\mathbb{F}_{p}$. The case for all $B^{i}_{\alpha,\beta},\alpha\in[v],\beta\in[m]$ follow similarly.
    Input: $\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{u\times v},\llbracket B\rrbracket_{d}\in\mathbb{F}_{p}^{v\times m}$.
    Output: $\llbracket C\rrbracket_{d}\in\mathbb{F}_{p}^{u\times m}$, where $C^{i}=A^{i}\cdot B^{i}$ and $\llbracket C_{\alpha,\beta}\rrbracket_{d}\in\mathbb{F}_{p}$ packs $k$ values $\boldsymbol{c}_{i},i\in[k]$ and $\boldsymbol{c}_{i}$ is inner product of the $\alpha$-th row of $A^{i}$ and the $\beta$-th column of $B^{i}$.
    Offline:
    1. 1.
      Generate $u\cdot m\cdot\ell$ random bits PSS $\llbracket R^{\prime\prime i}_{\alpha,\beta}\rrbracket_{d}\leftarrow\mathcal{F}_{RandomBits},\alpha\in[u],\beta\in[m],i\in[\ell]$, where the length of each $R^{\prime\prime i}_{\alpha,\beta}$ is $k$.
    2. 2.
      Get $\llbracket R_{\alpha,\beta}\rrbracket_{2d}=\sum_{i=0}^{\ell-1}2^{i}\cdot\llbracket R^{\prime\prime i}_{\alpha,\beta}\rrbracket_{d}\cdot\llbracket R^{\prime\prime i}_{\alpha,\beta}\rrbracket_{d}+\llbracket\boldsymbol{0}\rrbracket_{d}.$
    3. 3.
      Get $\llbracket R^{\prime}_{\alpha,\beta}\rrbracket_{d}=\sum_{i=\ell-\ell_{x}}^{\ell-1}2^{i-\ell_{x}}\cdot\llbracket R^{\prime\prime i}_{\alpha,\beta}\rrbracket_{d}.$
    Online:
    1. 1.
      Compute $\llbracket C\rrbracket_{2d}=\llbracket A\rrbracket_{d}\cdot\llbracket B\rrbracket_{d}+\llbracket R\rrbracket_{d}$ and send $\llbracket C\rrbracket_{2d}$ to the $P_{1}$.
    2. 2.
      $P_{1}$ reconstructs the matrix $C$, computes $C^{\prime}=C/2^{\ell_{x}}$ and distributes $\llbracket C^{\prime}\rrbracket_{d}$ to the other parties.
    3. 3.
      Compute $\llbracket C\rrbracket_{d}=\llbracket C^{\prime}\rrbracket_{d}-\llbracket R^{\prime}\rrbracket_{d}.$
    return $\llbracket C\rrbracket_{d}.$
    Algorithm 6 Parallel Matrix Mult with Trunc $\Pi_{PMatMult}^{Trunc}$
    Theorem 6.
    $\Pi_{PMatMult}^{Trunc}$ securely realizes $\mathcal{F}_{PMatMult}^{Trunc}$ in the $\mathcal{F}_{RandomBits}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    IV-B Inter-layer Connections
    The number of output channels in a convolutional layer is determined by the number of filters. Each filter produces an independent output channel. After accomplishing the convolution computation by using our parallel matrix multiplication protocol, the $k$ values that are selected respectively from the same position of $k$ output channels are packed into a single packed secret share as shown in Fig. 5. It should be noted that the results may require further processing before being propagated to the next layer.
    When the next layer performs a non-linear operation, the result can be directly utilized as input since non-linear layer operations are performed element-wise §V.
    When the next layer is an FC layer, seamless integration can be achieved without requiring any communication. The reason is that the results of the convolution computation require flattening to a vector and the flattened vector may be regarded as already packed through block-wise packing approach in §III. Correspondingly, we only need to pack the weight matrix according to the index of each element of the flattened vector.
    When the next layer still performs a Conv operation, a pack transformation is required for the result of the convolution computation since we need to pack $k$ copies of the input into a single packed secret share. To be specific, the pack transformation is that given a PSS $\llbracket\boldsymbol{x}\rrbracket_{d}$, it outputs $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}$, where $\boldsymbol{x}$ and $\boldsymbol{x}^{i}$ are two vectors of length $k,\boldsymbol{x}^{i}=(\boldsymbol{x}_{i},\cdots,\boldsymbol{x}_{i}),i\in[k]$. Fig. 5 presents an example. We formally define the functionality $\mathcal{F}_{PackTrans}$ in Appendix -A and provide a secure protocol implementation in Algorithm 7. In the offline, we generate $\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d},\llbracket\boldsymbol{r}\rrbracket_{d},$ where $\llbracket\boldsymbol{r}\rrbracket_{d}=(\boldsymbol{r}_{0},\cdots,\boldsymbol{r}_{k-1}),\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}=(\boldsymbol{r}_{i},\cdots,\boldsymbol{r}_{i}),i\in[k].$ In the online, parties compute $\llbracket\boldsymbol{z}\rrbracket_{d}=\llbracket\boldsymbol{x}\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{d}$ and reveal the vector $\boldsymbol{z}$ to $P_{1}$. Then, $P_{1}$ sets $\boldsymbol{z}^{i}=(\boldsymbol{z}_{i},\cdots,\boldsymbol{z}_{i})$ and distributes $\llbracket\boldsymbol{z}^{i}\rrbracket_{d}$. Finally, parties compute $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}=\llbracket\boldsymbol{z}^{i}\rrbracket_{d}-\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}$.
    To sum up, the online complexity is 1 round and $(n-1)(\frac{1}{2}+k)$ field elements per party. The offline complexity is 2 rounds and $(n-1)(\frac{2+k}{n-t}+\frac{2}{n})$ field elements per party.
    Figure 5: A toy example where $k=3$. Each pixel of the input tensor needs to be multiplied by each filter. $\boldsymbol{y}_{1},\boldsymbol{y}_{2},\boldsymbol{y}_{3}$ are packed into a single packed secret share. Three copies of $\boldsymbol{x}_{1}$ are packed into a single packed secret share. If the next layer is still a Conv operation, the result $\llbracket\boldsymbol{z}\rrbracket_{d}$ needs to be converted to $\llbracket\boldsymbol{z}^{i}\rrbracket_{d},i\in\{1,2,3\},\boldsymbol{z}=(\boldsymbol{z}_{1},\boldsymbol{z}_{2},\boldsymbol{z}_{3}),\boldsymbol{z}^{1}=(\boldsymbol{z}_{1},\boldsymbol{z}_{1},\boldsymbol{z}_{1}),\boldsymbol{z}^{2}=(\boldsymbol{z}_{2},\boldsymbol{z}_{2},\boldsymbol{z}_{2}),\boldsymbol{z}^{3}=(\boldsymbol{z}_{3},\boldsymbol{z}_{3},\boldsymbol{z}_{3})$ by pack transformation protocol. If the next layer is an FC operation, the result $\llbracket\boldsymbol{z}\rrbracket_{d}$ can become a flattened vector seamlessly and the weight matrix can be packed according to the index of each element of the flattened vector.
    Input: $\llbracket\boldsymbol{x}\rrbracket_{d}\in\mathbb{F}_{p}$.
    Output: $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}\in\mathbb{F}_{p},i\in[k]$, where $\boldsymbol{x}^{i}$ is a vector of length $k$ and each element is equivalent to $\boldsymbol{x}_{i}$.
    Offline:
    1. 1.
      Generare $\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}\leftarrow\mathcal{F}_{Random}$, where $\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}=(\boldsymbol{r}_{i},\cdots,\boldsymbol{r}_{i}),\boldsymbol{r}_{i}$ is a random value.
    2. 2.
      Compute $\llbracket\boldsymbol{r}\rrbracket_{d+k-1}=\sum_{i=0}^{k-1}\llbracket E_{i}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d},$ where $\boldsymbol{r}=(\boldsymbol{r}_{0},\cdots,\boldsymbol{r}_{k-1}).$
    3. 3.
      Get $\llbracket\boldsymbol{r}\rrbracket_{d}\leftarrow\mathcal{F}_{DegreeTrans}(\llbracket\boldsymbol{r}\rrbracket_{d+k-1}).$
    Online:
    1. 1.
      Compute $\llbracket\boldsymbol{z}\rrbracket_{d}=\llbracket\boldsymbol{x}\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{d}$ and send $\llbracket\boldsymbol{z}\rrbracket_{d}$ to the $P_{1}$.
    2. 2.
      $P_{1}$ reconstructs the vector $\boldsymbol{z}=(\boldsymbol{z}_{0},\cdots,\boldsymbol{z}_{k-1})$, sets $\boldsymbol{z}^{i}=(\boldsymbol{z}_{i},\boldsymbol{z}_{i},\cdots,\boldsymbol{z}_{i})$ and distributes $\llbracket\boldsymbol{z}^{i}\rrbracket_{d}$ to the other parties.
    3. 3.
      Compute $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}=\llbracket\boldsymbol{z}^{i}\rrbracket_{d}-\llbracket\boldsymbol{r^{\prime}}^{i}\rrbracket_{d},i\in[k].$
    return $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}.$
    Algorithm 7 Pack Transformation $\Pi_{PackTrans}$
    Theorem 7.
    $\Pi_{PackTrans}$ securely realizes $\mathcal{F}_{PackTrans}$ in the $(\mathcal{F}_{DegreeTrans},\mathcal{F}_{Random})$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    V Non-Linear Layers
    In this section, we introduce our efficient protocols for non-linear operations such as parallel prefix-ORs and bitwise less than in §V-A, parallel derivative of ReLU (DReLU) and ReLU in §V-B, and parallel Maxpool in §V-C.
    V-A Parallel Prefix-ORs and Bitwise Less Than
    We first describe our method to compute the prefix-ORs, which is the core and bottleneck to realizing the bitwise less-than functionality. The prefix-ORs operation ($b_{j}=\lor_{i=0}^{j}a_{i},a_{i}\in\{0,1\},j\in[\ell]$) can be implemented via its dual problem, that is, $\bar{b}_{j}=\land_{i=0}^{j}\bar{a}_{i}.$ It is worth noting that [5] designs a protocol for prefix-ORs with only one round of communication; however, [11] identified a privacy leakage issue in that protocol. In this work, we avoid this privacy leakage by utilizing functionality $\mathcal{F}_{Pmult-DN}$ and provide a secure protocol implementation for parallel prefix-ORs in Algorithm 8. $\mathcal{F}_{PreMult}$ denotes parallel prefix multiplication functionality and we formally define the functionality in Appendix -A. Further, $\mathcal{F}_{PreMult}$ can be securely implemented by directly employing $\mathcal{F}_{Pmult-DN}$ as a subroutine. As illustrated in Fig. 6, an 8-input prefix multiplication is depicted, where denotes a multiplication gate.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil$ rounds with $\lceil\ell\text{log}_{2}(\ell)\rceil\frac{n}{k}$ field elements per party. The offline communication complexity is 1 rounds with $\frac{n\ell}{k(n-t)}\lceil\text{log}_{2}(\ell)\rceil$ field elements per party.
    Input: $(\llbracket\boldsymbol{a}^{0}\rrbracket_{d},\cdots,\llbracket\boldsymbol{a}^{\ell-1}\rrbracket_{d})$, where $\boldsymbol{a}^{i}$ is a bit vector of length $k,\llbracket\boldsymbol{a}^{i}\rrbracket_{d}\in\mathbb{F}_{p},\boldsymbol{a}^{i}=(\boldsymbol{a}^{i}_{0},\cdots,\boldsymbol{a}^{i}_{k-1})$.
    Output: All prefix ORs $\llbracket\boldsymbol{a}^{0}\rrbracket_{d},\llbracket\boldsymbol{a}^{0}\lor\boldsymbol{a}^{1}\rrbracket_{d},\cdots,\llbracket\lor_{i=0}^{\ell-1}\boldsymbol{a}^{i}\rrbracket_{d}$, where $\lor_{i=0}^{\ell-1}\boldsymbol{a}^{i}=(\lor_{i=0}^{\ell-1}\boldsymbol{a}^{i}_{0},\cdots,\lor_{i=0}^{\ell-1}\boldsymbol{a}^{i}_{k-1}).$
    1. 1.
      Compute $\llbracket\boldsymbol{b}^{i}\rrbracket_{d}=1-\llbracket\boldsymbol{a}^{i}\rrbracket_{d}$.
    2. 2.
      Get $\llbracket\boldsymbol{c}^{i}\rrbracket_{d}\leftarrow\mathcal{F}_{PreMult}(\llbracket\boldsymbol{b}^{0}\rrbracket_{d},\cdots,\llbracket\boldsymbol{b}^{\ell-1}\rrbracket_{d})$.
    3. 3.
      Compute $1-\llbracket\boldsymbol{c}^{0}\rrbracket_{d},\cdots,1-\llbracket\boldsymbol{c}^{\ell-1}\rrbracket_{d}$.
    return $\llbracket\boldsymbol{a}^{0}\rrbracket_{d},\llbracket\boldsymbol{a}^{0}\lor\boldsymbol{a}^{1}\rrbracket_{d},\cdots,\llbracket\lor_{i=0}^{\ell-1}\boldsymbol{a}^{i}\rrbracket_{d}.$
    Algorithm 8 Prefix-ORs $\Pi_{PreOR}$
    Theorem 8.
    $\Pi_{PreOR}$ securely realizes $\mathcal{F}_{PreOR}$ in the $\mathcal{F}_{PreMult}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Based on PSS, multiple bitwise less-than circuits can be securely computed in parallel. Specifically, given $k$ public values $a_{0},\cdots,a_{k-1}$ and $\ell$ shared bits vectors $\llbracket\boldsymbol{b}^{i}\rrbracket_{d},i\in[\ell],\boldsymbol{b}^{i}\in\{0,\cdots,1\}^{k}$, it outputs the share of $(a_{j}<b_{j}),b_{j}=\sum_{i=0}^{\ell-1}2^{i}\boldsymbol{b}^{i}_{j},j\in[k]$. Our protocol $\Pi_{Bitwise-LT}$ follows the method in [5], with the substitution of our secure prefix-ORs subprotocol $\Pi_{PreOR}$. Algorithm 9 depicts the protocol $\Pi_{Bitwise-LT}$. In Step 4, the XOR operation can be computed by $\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\oplus\llbracket\boldsymbol{b}^{i}\rrbracket_{d}=\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}+\llbracket\boldsymbol{b}^{i}\rrbracket_{d}-2\cdot\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{b}^{i}\rrbracket_{d}.$ We formally define the functionality $\mathcal{F}_{Xor}$ in Appendix -A and provide a secure protocol implementation in Appendix -B.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+2$ rounds with $\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil)$ field elements per party. The offline communication complexity is 1 rounds with $\frac{n}{k(n-t)}(4+\lceil\ell\text{log}_{2}(\ell)\rceil)$ field elements per party.
    Theorem 9.
    $\Pi_{Bitwise-LT}$ securely realizes $\mathcal{F}_{Bitwise-LT}$ in the $(\mathcal{F}_{PreOR},\mathcal{F}_{DegreeTrans},\mathcal{F}_{Xor})$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Figure 6: An 8-input prefix multiplication. denotes a multiplication gate that can be securely implemented by $\mathcal{F}_{Pmult-DN}$.
    Input: $k$ public values $a_{0},\cdots,a_{k-1}$, $\ell$ shared bits vectors $\llbracket\boldsymbol{b}^{i}\rrbracket_{d},i\in[\ell],\boldsymbol{b}^{i}\in\{0,\cdots,1\}^{k},b_{j}=\sum_{i=0}^{\ell-1}2^{i}\boldsymbol{b}^{i}_{j}$.
    Output: $\llbracket a_{j}<b_{j}\rrbracket_{d},j\in[k]$
    1. 1.
      Let $a_{0,0},\cdots,a_{0,\ell-1},\cdots,a_{k-1,0},\cdots,a_{k-1,\ell-1}$ be the decomposed bits of $a_{0},\cdots,a_{k-1}$.
    2. 2.
      Set $\bar{a}_{j,i}=1-a_{j,i}$ and $\llbracket\bar{\boldsymbol{b}}^{i}\rrbracket_{d}=1-\llbracket\boldsymbol{b}^{i}\rrbracket_{d},i\in[\ell],j\in[k].$
    3. 3.
      Compute $\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}$, where $\boldsymbol{a}^{i}=(\bar{a}_{0,i},\cdots,\bar{a}_{k,i}),i\in[\ell]$.
    4. 4.
      Compute $\llbracket\boldsymbol{c}^{i}\rrbracket_{d}\leftarrow\mathcal{F}_{Xor}(\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1},\llbracket\bar{\boldsymbol{b}}^{i}\rrbracket_{d})$.
    5. 5.
      Get $\llbracket\boldsymbol{f}^{\ell-1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{f}^{0}\rrbracket_{d}\leftarrow\mathcal{F}_{PreOR}(\llbracket\boldsymbol{c}^{\ell-1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{c}^{0}\rrbracket_{d})$.
    6. 6.
      Set $\llbracket\boldsymbol{h}^{i}\rrbracket_{d}=\llbracket\boldsymbol{f}^{i}\rrbracket_{d}-\llbracket\boldsymbol{f}^{i+1}\rrbracket_{d}$, where $\llbracket\boldsymbol{h}^{\ell-1}\rrbracket_{d}=\llbracket\boldsymbol{f}^{\ell-1}\rrbracket_{d}.$
    7. 7.
      Compute $\llbracket a_{j}<b_{j}\rrbracket_{d+k-1}=\sum_{i=0}^{\ell-1}\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{h}^{i}\rrbracket_{d}.$
    8. 8.
      Output $\llbracket a_{j}<b_{j}\rrbracket_{d}\leftarrow\mathcal{F}_{DegreeTrans}(\llbracket a_{j}<b_{j}\rrbracket_{d+k-1}),j\in[k].$
    return $\llbracket a_{j}<b_{j}\rrbracket_{d},j\in[k]$.
    Algorithm 9 Bitwise Less Than $\Pi_{Bitwise-LT}$
    V-B Parallel DReLU and ReLU
    ReLU is a common activation function in neural networks. It can be represented as $\text{ReLU}(x)=\text{Max}(0,x)$ and the derivative of ReLU (DReLU) is 1 for $x\geq 0$ and 0 for $x<0$.
[…]
    Since integers are represented in binary two’s complement form, the most significant bit of an integer is related to DReLU. Furthermore, Mersenne primes are necessarily odd, which implies that $\text{MSB}(a)=\text{LSB}(2a)$ according to [12]. Therefore, it follows that $\text{DReLU}(a)=1-\text{LSB}(2a)$. As in [12, 5], in order to compute securely $\text{LSB}(2a)$, we mask $2a$ with a random value $r$ such that $y=2a+r$ and reveal $y$. Consequently, $\text{LSB}(y)=\text{LSB}(2a)\oplus\text{LSB}(r)$ if $y\geq r$, and $\text{LSB}(y)=\text{LSB}(2a)\oplus\text{LSB}(r)\oplus 1$ otherwise and we can get $\text{LSB}(2a)=\text{LSB}(y)\oplus\text{LSB}(r)\oplus(y<r)$.
    We employ the PSS to enable parallel computation of multiple DReLU. Specifically, given $k$ public values $a_{0},\cdots,a_{k-1}$, it outputs the of $\text{DReLU}(a_{i}),i\in[k]$. We provide a secure protocol implementation in Algorithm 10.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+5$ rounds with $\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil+4)+n$ field elements per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(13+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)$ field elements per party.
    Note that $\text{ReLU}(a)=\text{DReLU}(a)\cdot a$. So we can easily implement parallel computation of multiple ReLU functions. The secure protocol is in Algorithm 11.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+6$ rounds with $\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil+6)+n$ field elements per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(15+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)$ field elements per party.
    Input: $\llbracket\boldsymbol{a}\rrbracket_{d},\boldsymbol{a}$ is a vector with length $k$.
    Output: $\llbracket\text{DReLU}(\boldsymbol{a})\rrbracket_{d}.$
    Offline:
    1. 1.
      Generate $\ell$ packed bit secret shares $\llbracket\boldsymbol{r}^{i}\rrbracket_{d}\leftarrow\mathcal{F}_{RandomBits},i\in[\ell]$, where $\boldsymbol{r}^{i}$ is a bit vector with length $k$.
    2. 2.
      Compute $\llbracket\boldsymbol{r}\rrbracket_{d}=\sum_{i=0}^{\ell-1}2^{i}\cdot\llbracket\boldsymbol{r}^{i}\rrbracket_{d},$ where $\boldsymbol{r}_{j}=\sum_{i=0}^{\ell-1}2^{i}\cdot\boldsymbol{r}^{i}_{j},j\in[k].$
    Online:
    1. 1.
      Compute $\llbracket\boldsymbol{y}\rrbracket_{d}=2\llbracket\boldsymbol{a}\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{d}$, send $\llbracket\boldsymbol{y}\rrbracket_{d}$ to $P_{1}$ and $P_{1}$ reconstructs and opens $\boldsymbol{y}=(\boldsymbol{y_{0}},\cdots,\boldsymbol{y}_{k-1}).$
    2. 2.
      Compute Let $\boldsymbol{y}_{0,0},\cdots,\boldsymbol{y}_{0,\ell-1},\cdots,\boldsymbol{y}_{k-1,0},\cdots,\boldsymbol{y}_{k-1,\ell-1}$ be the decomposed bits of $\boldsymbol{y}_{0},\cdots,\boldsymbol{y}_{k-1}$.
    3. 3.
      Compute $\llbracket\boldsymbol{y}^{0}\rrbracket_{k-1}$, where $\boldsymbol{y}^{0}=(\boldsymbol{y}_{0,0},\cdots,\boldsymbol{y}_{k,0})$.
    4. 4.
      Compute $\llbracket\boldsymbol{b}\rrbracket_{d}\leftarrow\mathcal{F}_{Xor}(\llbracket\boldsymbol{y}^{0}\rrbracket_{k-1},\llbracket\boldsymbol{r}^{0}\rrbracket_{d})$.
    5. 5.
      Get $\llbracket\boldsymbol{c}\rrbracket_{d}\leftarrow\mathcal{F}_{Bitwise-LT}(\boldsymbol{y}_{0},\cdots,\boldsymbol{y}_{k-1},\llbracket\boldsymbol{r}^{i}\rrbracket_{d}),i\in[\ell]$.
    6. 6.
      Compute $\llbracket\boldsymbol{f}\rrbracket_{d}\leftarrow\mathcal{F}_{Xor}(\llbracket\boldsymbol{b}\rrbracket_{d},\llbracket\boldsymbol{c}\rrbracket_{d})$.
    7. 7.
      Output $1-\llbracket\boldsymbol{f}\rrbracket_{d}.$
    return $\llbracket\text{DReLU}(\boldsymbol{a})\rrbracket_{d}.$
    Algorithm 10 DReLU $\Pi_{DReLU}$
    Theorem 10.
    $\Pi_{DReLU}$ securely realizes $\mathcal{F}_{DReLU}$ in the $(\mathcal{F}_{Bitwise-LT},\mathcal{F}_{Xor})$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Input: $\llbracket\boldsymbol{a}\rrbracket_{d},\boldsymbol{a}$ is a vector with length $k$.
    Output: $\llbracket\text{ReLU}(\boldsymbol{a})\rrbracket_{d}.$
    1. 1.
      Get $\llbracket\boldsymbol{b}\rrbracket_{d}\leftarrow\mathcal{F}_{DReLU}(\boldsymbol{a}).$
    2. 2.
      Output $\llbracket\text{ReLU}(\boldsymbol{a})\rrbracket_{d}\leftarrow\mathcal{F}_{Pmult-DN}(\llbracket\boldsymbol{a}\rrbracket_{d},\llbracket\boldsymbol{b}\rrbracket_{d}).$
    return $\llbracket\text{ReLU}(\boldsymbol{a})\rrbracket_{d}.$
    Algorithm 11 ReLU $\Pi_{ReLU}$
    Theorem 11.
    $\Pi_{ReLU}$ securely realizes $\mathcal{F}_{ReLU}$ in the $(\mathcal{F}_{DReLU},\mathcal{F}_{PMult-DN})$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    V-C Parallel Maxpool
    As described in §II-C, Maxpool is the key component in convolutional neural networks (CNN). The Maxpool can be reformulated as finding the maximum value in a vector of length $m$. We can represent $\text{Max}(a,b)$ as follows:
    $\text{Max}(a,b)=\text{ReLU}(a-b)+b$
    So Maxpool can be realized through ReLU. Assume that $m$ is a power of two. To identify the maximum among $m$ elements, the values are first grouped into pairs, and a comparison is performed within each pair to determine the local maximum, yielding a new vector of size $m/2$. This procedure is applied recursively over $\text{log}_{2}(m)$ iterations until a single maximum value is obtained.
    Similarly, we use PSS to enable the parallel computation of multiple Maxpool operations. Specifically, given $m$ shared vectors $\llbracket\boldsymbol{a}^{i}\rrbracket_{d},$ where $\boldsymbol{a}^{i}$ is vector with length $k,i\in[m]$. It outputs the share of $\text{Max}(\boldsymbol{a}^{1}_{j},\cdots,\boldsymbol{a}^{m}_{j}),j\in[k]$. We provide a secure protocol $\Pi_{Maxpool}$ implementation in Algorithm 12.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(m)\rceil(\lceil\text{log}_{2}(\ell)\rceil+6)$ rounds with $(\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil+6)+n)(m-1)$ field elements per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(15+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)(m-1)$ field elements per party.
    TABLE I: Communication costs of per party of Liu et al. [5] and our protocol over different datasets and networks for secure inference. Communication is in MB. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
    n t protocol k MiniONN(MNIST) LeNet(MNIST) AlexNet(CIFAR10) VGG16(CIFAR10)
    offline online total offline online total offline online total offline online total
    11 3  [5] - 71.77 40.45 112.22 105.63 59.53 165.16 165.04 99.40 264.44 1975.53 1113.65 3089.18
    Ours 3 46.36 21.42 67.78 63.28 29.24 92.52 89.54 46.92 136.46 1157.48 534.87 1692.35
    4  [5] - 67.81 43.17 110.98 99.81 63.54 163.35 154.42 106.07 260.49 1866.63 1188.51 3055.14
    Ours 2 55.78 26.59 82.37 82.10 39.14 121.24 121.60 65.41 187.01 1537.66 732.92 2270.58
    21 3  [5] - 78.73 36.97 115.70 115.89 54.42 170.31 183.03 90.86 273.89 2167.38 1017.94 3185.32
    Ours 8 22.63 8.32 30.95 39.20 14.40 53.60 47.93 20.44 68.37 624.24 229.78 854.02
    7  [5] - 72.29 42.93 115.22 106.40 63.20 169.60 169.30 105.50 270.80 1989.81 1182.10 3171.91
    Ours 4 34.02 14.49 48.51 50.51 21.50 72.01 73.15 35.63 108.78 938.21 399.58 1337.79
    31 3  [5] - 80.89 35.62 116.51 119.07 52.42 171.49 188.64 87.54 276.18 2226.84 980.78 3207.62
    Ours 13 29.13 8.74 37.87 32.80 9.82 42.62 40.32 14.23 54.55 505.03 151.77 656.80
    12  [5] - 70.30 44.85 115.15 103.47 66.02 169.49 159.74 110.21 269.95 OOM OOM OOM
    Ours 4 33.95 14.20 48.15 50.39 21.08 71.47 72.84 34.92 107.76 936.08 391.66 1327.74
    63 3  [5] - 83.06 34.11 117.17 122.26 50.21 172.47 194.29 83.85 278.14 OOM OOM OOM
    Ours 29 24.86 4.68 29.54 28.04 5.25 33.29 33.23 7.51 40.74 458.81 86.83 545.64
    20  [5] - 75.26 42.84 118.10 110.78 63.05 173.83 172.48 105.27 277.75 OOM OOM OOM
    Ours 12 27.44 7.70 35.14 32.27 9.05 41.32 38.49 12.71 51.20 542.71 152.61 695.32
    Input: $\llbracket\boldsymbol{a}^{i}\rrbracket_{d},\boldsymbol{a}$ is a vector with length $k,i\in[m]$. Assuming $m$ is a power of two.
    Output: $\llbracket\boldsymbol{b}\rrbracket_{d},$ where $\boldsymbol{b}$ is a vector with length $k$ and $\boldsymbol{b}_{j}=\text{Max}(\boldsymbol{a}^{1}_{j},\cdots,\boldsymbol{a}^{m}_{j}),j\in[k].$
    1. 1.
      If $m=1:$ Output $\llbracket\boldsymbol{a}^{0}\rrbracket_{d}$.
    2. 2.
      Set $s=m$ and $\boldsymbol{c^{1}}=(\llbracket\boldsymbol{a}^{0}\rrbracket_{d},\cdots,\llbracket\boldsymbol{a}^{m-1}\rrbracket_{d}).$
    3. 3.
      for $i\leftarrow 0$ to $\text{log}_{2}(m)-1$ do
       Set $s=s/2$, $(\llbracket\boldsymbol{f}^{0}\rrbracket_{d},\cdots,\llbracket\boldsymbol{f}^{s-1}\rrbracket_{d},\llbracket\boldsymbol{h}^{0}\rrbracket_{d},\cdots,\llbracket\boldsymbol{h}^{s-1}\rrbracket_{d})=\boldsymbol{c}^{i}.$ for $j\leftarrow 0$ to s1s-1 do
        $\llbracket\boldsymbol{g}^{j}\rrbracket_{d}\leftarrow\mathcal{F}_{ReLU}(\llbracket\boldsymbol{f}^{j}-\boldsymbol{h}^{j}\rrbracket_{d})+\boldsymbol{h}^{j}$
      If $s=1:$ Output $\llbracket\boldsymbol{g}^{0}\rrbracket_{d}$. Else: Set $\boldsymbol{c}^{i+1}\leftarrow(\llbracket\boldsymbol{g}^{0}\rrbracket_{d},\cdots,\llbracket\boldsymbol{g}^{s-1}\rrbracket_{d})$.
    return $\llbracket\boldsymbol{b}\rrbracket_{d}.$
    Algorithm 12 Maxpool $\Pi_{Maxpool}$
    Theorem 12.
    $\Pi_{Maxpool}$ securely realizes $\mathcal{F}_{Maxpool}$ in the $\mathcal{F}_{ReLU}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    VI Evaluation
    TABLE II: Running time of Liu et al. [5] and our protocol over different datasets and networks for secure inference in WAN. Running time is in seconds. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
    n t protocol k MiniONN(MNIST) LeNet(MNIST) AlexNet(CIFAR10) VGG16(CIFAR10)
    offline online total offline online total offline online total offline online total
    11 3  [5] - 17.30 9.64 26.94 23.80 13.92 37.72 35.35 22.15 57.50 426.93 182.71 609.64
    Ours 3 18.63 13.52 32.15 24.04 17.06 41.10 34.93 30.86 65.79 396.67 261.88 658.55
    4  [5] - 16.82 9.93 26.75 23.51 13.21 36.72 34.34 22.43 56.77 420.59 183.26 603.85
    Ours 2 19.11 15.21 34.32 27.50 21.76 49.26 39.24 37.38 76.62 451.03 335.75 786.78
    21 3  [5] - 32.82 15.68 48.50 48.14 22.02 70.16 64.87 35.80 100.67 914.20 307.68 1221.88
    Ours 8 23.08 11.84 34.92 39.41 18.66 58.07 46.28 29.24 75.52 574.19 239.57 813.76
    7  [5] - 30.16 15.72 45.88 44.11 22.05 66.16 62.81 36.66 99.47 910.92 357.59 1268.51
    Ours 4 27.76 16.93 44.69 39.87 23.74 63.61 53.99 40.58 94.57 684.68 360.72 1045.40
    31 3  [5] - 48.14 20.09 68.23 67.59 28.47 96.06 95.06 46.27 141.33 1203.35 461.91 1665.26
    Ours 13 49.91 17.84 67.75 55.79 19.59 75.38 65.68 30.40 96.08 793.62 240.29 1033.91
    12  [5] - 44.12 22.01 66.13 63.73 30.32 94.05 90.15 49.95 140.10 OOM OOM OOM
    Ours 4 39.76 22.66 62.42 57.29 31.31 88.60 77.38 52.89 130.27 1029.76 498.14 1527.90
    63 3  [5] - 89.99 36.69 126.68 131.76 55.08 186.84 188.09 86.27 274.36 OOM OOM OOM
    Ours 29 102.48 20.09 122.57 112.61 22.09 134.70 129.51 33.07 162.58 1738.20 343.99 2082.19
    20  [5] - 87.73 45.19 132.92 124.20 64.03 188.23 180.91 106.59 287.50 OOM OOM OOM
    Ours 12 89.40 26.46 115.86 103.31 30.16 133.47 119.13 44.83 163.96 1791.95 476.84 2268.79
    TABLE III: Running time of Liu et al. [5] and our protocol over different datasets and networks for secure inference in LAN. Running time is in seconds. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
    n t protocol k MiniONN(MNIST) LeNet(MNIST) AlexNet(CIFAR10) VGG16(CIFAR10)
    offline online total offline online total offline online total offline online total
    11 3  [5] - 1.13 0.20 1.33 1.70 0.28 1.98 2.74 0.46 3.20 39.30 5.12 44.42
    Ours 3 0.93 0.33 1.26 1.29 0.46 1.75 1.77 0.74 2.51 29.26 8.26 37.52
    4  [5] - 1.48 0.21 1.69 2.22 0.31 2.53 3.48 0.51 3.99 48.96 5.83 54.79
    Ours 2 1.27 0.47 1.74 1.75 0.68 2.43 2.66 1.17 3.83 42.37 13.06 55.43
    21 3  [5] - 1.17 0.23 1.40 1.69 0.29 1.98 2.79 0.57 3.36 23.22 6.14 29.36
    Ours 8 0.74 0.34 1.08 1.37 0.60 1.97 1.54 0.85 2.39 21.33 9.83 31.16
    7  [5] - 1.93 0.28 2.21 2.84 0.44 3.28 4.78 0.69 5.47 38.62 8.68 47.30
    Ours 4 1.10 0.45 1.55 1.71 0.72 2.43 2.21 1.19 3.40 35.01 13.25 48.26
    31 3  [5] - 1.15 0.25 1.40 1.76 0.36 2.12 2.68 0.89 3.57 38.25 7.09 45.34
    Ours 13 1.37 0.68 2.05 1.59 0.79 2.35 1.87 1.15 3.02 25.83 13.63 39.46
    12  [5] - 3.15 0.44 3.59 4.79 0.64 5.43 7.54 1.35 8.89 OOM OOM OOM
    Ours 4 1.56 0.68 2.24 2.34 1.24 3.58 3.28 1.62 4.90 51.42 21.59 73.01
    63 3  [5] - 1.77 0.86 2.63 2.76 0.51 3.27 3.91 1.97 5.88 OOM OOM OOM
    Ours 29 2.65 1.72 4.37 3.06 1.93 4.99 3.52 2.78 6.30 50.31 33.26 83.57
    20  [5] - 4.68 1.16 5.84 7.27 1.13 8.40 12.59 2.61 15.20 OOM OOM OOM
    Ours 12 2.42 1.50 3.92 2.83 1.85 4.68 3.35 2.48 5.83 82.09 57.29 139.38
    In this section, we evaluate the performance of secure inference on 4 networks using MNIST and CIFAR10, and analyse the experimental results.
    VI-A Experiments Environment
    We implement this framework on top of Liu et al. [5] in C++ [^1] The source code will be open-sourced soon in github. and run our experiments on 11 Alibaba Cloud servers, each of which was equipped with an AMD EPYC 32-core processor with 64GB of RAM. To simulate different network conditions, we use Linux tc tool to simulate a wide-area network (WAN, RTT: 40 ms, 100 Mbps), and a local-area network (LAN, RTT: 0.2 ms, 10 Gbps). RTT denotes Round-Trip Time. We completed performance evaluations for inference in various settings involving different numbers of parties, including 11-party computation (11PC), 21PC, 31PC, and 63PC. In addition, we rerun the protocol of Liu et al. [5] under the same adversarial model. Since the single-round communication prefix multiplication protocol in  [5] suffers from privacy leakage issues, we replace it with a logarithmic-round prefix multiplication protocol based on the DN multiplication protocol. In the experiment, all arithmetic operations on the shares are performed modulo $p=2^{61}-1$, with a fixed-point precision of 13 bits. All evaluations are executed 10 times, and record the average.
    VI-B Neural Networks & Datasets
    We conducted evaluation on 4 standard neural networks: MiniONN [18] from the secure inference community, and LeNet [36], AlexNet [38], and VGG16 [39] from the deep learning community. We select 2 standard benchmarking datasets: MNIST [36] and CIFAR-10 [37].
    VI-C Communication Costs of Secure Inference
    The communication costs of the protocol from Liu et al. [5] and our protocol on various datasets and neural networks are shown in Table I. Compared to Liu et al. [5], we achieve a significant advantage in terms of offline communication, online communication, and total communication costs. Moreover, our analysis reveals a progressively more pronounced advantage along the diagonal from the upper-left to the lower-right corner of the table. For instance, compared to Liu et al. [5], our approach improves the communication by upto $5.85\times$, $11.17\times$, and $6.83\times$ in offline, online, and total communication overhead respectively for AlexNet. Additionally, for VGG16, Liu et al. [5] suffers from memory overflow under both the 31PC where 12 parties are controlled by the adversary and 63PC, whereas our approach continues to execute successfully. This advantage primarily stems from the fact that more parties enable the selection of larger values for $k$ in PSS, where increased $k$ confers greater benefits for deeper neural networks and larger datasets.
    VI-D Running Time of Secure Inference
    VI-D1 WAN
    The running time of Liu et al. [5] and ours in WAN is shown in Table II. Similar to Table I, our advantages become increasingly significant from the top-left to the bottom-right of the table. However, it is worth noting that in the case of 11PC, there is no improvement compared to Liu et al. [5]. This phenomenon primarily stems from our scheme is significantly affected by computational overhead, and when a small $k$ is selected, the communication efficiency improvement is limited. Furthermore, we can find that when the communication overhead increases by more than $3.80\times$, our online running time will be significantly improved, with a maximum increase of $2.61\times$. At the same time, in the case of more parties and a deeper neural network, the offline running time and the total running time can be increased by up to $1.52\times$ and $1.75\times$ respectively.
    VI-D2 LAN
    The running time of Liu et al. [5] and ours in LAN is shown in Table III. We observe that in the online phase, our running time shows little to no advantage compared to Liu et al. [5], as our scheme is more heavily affected by computational overhead. However, the offline and the total running time, as the number of parties increases and the network deepens, will also have significant improvements, with the maximum increase being $3.76\times$ and $2.61\times$ respectively.
    In a word, our improvements in the LAN setting are less pronounced compared to those in the WAN setting. The reason lies in the higher computational overhead required by our scheme. Fortunately, as the number of parties increases and the neural networks become deeper, our scheme exhibits more substantial advantages, mainly owing to our efficient vector-matrix multiplication, parallel convolution computation, and parallel non-linear computation based on PSS. Moreover, MPC protocols are more commonly deployed in WAN settings, so our improvements hold practical significance.
    VII Related Work
    Secure neural network inference has been a prominent research focus in recent years. Based on the number of participating parties, existing works can be broadly categorized into two categories: those involving a small number of parties (e.g., 2PC, 3PC, 4PC) and those designed for a large number of parties (nnPC).
    In the area of two-party computation (2PC), ABY [20] and its follow-up work ABY2.0 [21] design a framework that combines secure computation schemes based on arithmetic sharing, boolean sharing, and Yao’s garbled circuits (GC). SecureML [17] proposes a scheme in the two-server outsourced computation models with secret sharing and GC. Meanwhile, MiniONN [18] designs fast matrix multiplication protocols. Gazelle [6] and Delphi [19] combine techniques from HE and MPC to achieve fast private inference. CrypTFlow2 [24] designs an efficient comparison protocol, and its followup works [41, 42, 40] focus on improving performance. In addition, there exists many works [7, 44, 43] for large language model inference combines techniques from HE and secret sharing. In order to solve the performance bottleneck of 2PC, Chameleon [22] employed a semi-honest third-party to generate correlated randomness for multiplication triplets in offline. Further, for improving performance in non-linear functions such as ReLU, SecureNN [12] constructs novel protocols for comparsion with the help of a third-party. For providing better overall efficiency, some works ABY3 [9], Falcon[27], Swift [30] are propoed based on 3PC replicated secret sharing. Additionally,  [45, 28] and [34, 31, 33, 32] focus on resisting malicious adversaries and require much more costs in 3PC and 4PC, respectively.
    Instead, a critical limitation common to all these representative works lies in their restriction to a maximum of four participating parties, consequently exhibiting inadequate scalability. In order to solve the bottleneck, many works are proposed. Motion [47] designs an nnPC framework that combines secure computation schemes based on arithmetic sharing, boolean sharing, and GC. Howerver, since computation is required between every pair of parties in Motion, the overall efficiency is limited. Alessandro et al. [46] proposes a nnPC framework by utilizing replicated secret sharing, but its sharing structure is relatively complex and incurs high communication overhead for multiplication operations. Liu et al. [5] first constructs a practical inference framework using Shamir secret sharing. Unfortunately, when dealing with deep neural networks such as VGG16, the communication overhead remains substantial. Franklin et al. [1] introduced the PSS and its follow-up works [8, 3, 48] focus on improving multiplication operation and resisting malicious adversaries in theory. In this work, we mainly achieve efficient vector-matrix multiplication, convolution computation and non-linear operations for more practical secure inference using PSS.
    VIII Conclusion & Future Work
    In this paper, we construct secure blocks based on packed Shamir secret sharing for high-throughput and scalable secure inference. Extensive evaluations also present our improvements. For future work, we are willing to extend to malicious adversary scenarios and improve computational efficiency by GPU.
    -A Functionality Description
    Functionality $\mathcal{F}_{Random}$ Input: The functionality receives no inputs. Output: Compute the following 1. Randomly sample a vector $\boldsymbol{r}\in\mathbb{F}_{p}^{k}$. 2. Distribute shares of $\llbracket\boldsymbol{r}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{RandomBits}$ Input: The functionality receives no inputs. Output: Compute the following 1. Randomly sample a bit string $\boldsymbol{r}\in\mathbb{F}_{2}^{k}$. 2. Distribute shares of $\llbracket\boldsymbol{r}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{DegreeTrans}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}$ (resp. $a|_{s}]_{d_{2}}$), where $d\leq d_{1}\leq n-1,d-k+1\leq d_{2}\leq n-1$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{k}$ (resp. $a\in\mathbb{F}_{p}$). 2. Distribute shares of $\llbracket\boldsymbol{a}\rrbracket_{d_{3}}$ (resp. $a|_{s}]_{d_{4}}$) to the parties, where $d\leq d_{3}\leq n-1,d-k+1\leq d_{4}\leq n-1,d_{3}\neq d_{1},d_{4}\neq d_{2}.$
    Functionality $\mathcal{F}_{Pmult-DN}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}\rrbracket_{d}$ and $\llbracket\boldsymbol{b}\rrbracket_{d}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{k}$ and $\boldsymbol{b}\in\mathbb{F}_{p}^{k}$ to compute $\boldsymbol{c}_{i}=\boldsymbol{a}_{i}\boldsymbol{b}_{i},i\in[k]$. 2. Distribute shares of $\llbracket\boldsymbol{c}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{PMatMult}^{Trunc}$ Input: The functionality receives inputs $\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{u\times v}$ and $\llbracket B\rrbracket_{d}\in\mathbb{F}_{p}^{v\times w}$. Output: Compute the following 1. Reconstruct $A^{i}\in\mathbb{F}_{p}^{u\times v}$ and $B^{i}\in\mathbb{F}_{p}^{v\times w},i\in[k]$ to compute $C^{\prime i}=A^{i}\cdot B^{i},i\in[k]$. 2. Compute $C^{i}=C^{\prime i}_{\alpha,\beta}/2^{\ell_{x}},i\in[k],\alpha\in[u],\beta\in[w]$. 3. Distribute shares of $\llbracket C\rrbracket_{d}\in\mathbb{F}_{p}^{u\times w}$ to the parties.
    Functionality $\mathcal{F}_{VecMatMult}$ Input: The functionality receives inputs a vector $\llbracket\boldsymbol{a}\rrbracket_{d}\in\mathbb{F}_{q}^{u}$ and a matrix $\llbracket\boldsymbol{A}\rrbracket_{d}\in\mathbb{F}_{q}^{u\times vk}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{ku}$ and $\boldsymbol{A}\in\mathbb{F}_{p}^{ku\times vk}$ to compute $\boldsymbol{c}=\boldsymbol{a}\cdot A$. 2. Distribute shares of $\llbracket\boldsymbol{c}\rrbracket_{d}\in\mathbb{F}_{q}^{v}$ to the parties.
    Functionality $\mathcal{F}_{VecMatMult}^{Trunc}$ Input: The functionality receives inputs a vector $\llbracket\boldsymbol{a}\rrbracket_{d}\in\mathbb{F}_{q}^{u}$ and a matrix $\llbracket\boldsymbol{A}\rrbracket_{d}\in\mathbb{F}_{q}^{u\times vk}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{ku}$ and $\boldsymbol{A}\in\mathbb{F}_{p}^{ku\times vk}.$ 2. Compute $\boldsymbol{c}=\boldsymbol{a}\cdot A$ and compute $\boldsymbol{c}_{i}^{\prime}=\boldsymbol{c}_{i}/2^{\ell_{x}}$. 3. Distribute shares of $\llbracket\boldsymbol{c}^{\prime}\rrbracket_{d}\in\mathbb{F}_{q}^{v}$ to the parties.
    Functionality $\mathcal{F}_{RandomPairs}$ Input: The functionality receives no inputs. Output: Compute the following 1. Randomly sample a vecotr $\boldsymbol{r}\in\mathbb{F}_{p}^{kv}$ and compute a vector $\boldsymbol{r}^{\prime}=(\sum_{i=0}^{k-1}\boldsymbol{r}_{i},\cdots,\sum_{i=(v-1)k}^{kv-1}\boldsymbol{r}_{i})\in\mathbb{F}_{p}^{v}$. 2. Distribute shares of the random pairs $(\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d})$ to the parties.
    Functionality $\mathcal{F}_{TruncTriple}$ Input: The functionality receives no inputs. Output: Compute the following 1. Randomly sample a vecotr $\boldsymbol{r}\in\mathbb{F}_{p}^{kv}$ and compute a vector $\boldsymbol{r}^{\prime}=(\frac{\sum_{i=0}^{k-1}\boldsymbol{r}_{i}}{2^{\ell_{x}}},\cdots,\frac{\sum_{i=(v-1)k}^{kv-1}\boldsymbol{r}_{i}}{2^{\ell_{x}}})\in\mathbb{F}_{p}^{v}$. 2. Distribute shares of the random pairs $(\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d})$ to the parties.
    Functionality $\mathcal{F}_{PackTrans}$ Input: The functionality receives inputs $\llbracket\boldsymbol{x}\rrbracket_{d}.$ Output: Compute the following 1. Reconstruct $\boldsymbol{x}=(\boldsymbol{x}_{0},\cdots,\boldsymbol{x}_{k-1})$ and set $\boldsymbol{x}^{i}=(\boldsymbol{x_{i}},\cdots,\boldsymbol{x_{i}})\in\mathbb{F}_{p}^{k},i\in[k]$. 2. Distribute shares of $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{PreMult}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}^{0}\rrbracket_{d},\cdots,$ $\llbracket\boldsymbol{a}^{\ell-1}\rrbracket_{d},$ where $\boldsymbol{a}^{i}$ is a vector of length $k$, $\boldsymbol{a}^{i}=(\boldsymbol{a}_{0}^{i},\cdots,\boldsymbol{a}_{k-1}^{i}).$ Output: Compute the following 1. Reconstruct $(\boldsymbol{a}^{0}_{0},\cdots,\boldsymbol{a}^{0}_{k-1},\boldsymbol{a}^{1}_{0},\cdots,\boldsymbol{a}^{1}_{k-1},\cdots,\boldsymbol{a}^{\ell-1}_{0},\cdots,\boldsymbol{a}^{\ell-1}_{k-1})$. 2. Compute all prefix products $\boldsymbol{a}^{\prime 0}=(\boldsymbol{a}_{0}^{0},\cdots,\boldsymbol{a}_{k-1}^{0}),$ $\boldsymbol{a}^{\prime 1}=(\boldsymbol{a}_{0}^{0}\cdot\boldsymbol{a}_{0}^{1},\cdots,\boldsymbol{a}_{k-1}^{0}\cdot\boldsymbol{a}_{k-1}^{1})$, $\boldsymbol{a}^{\prime\ell-1}=(\Pi_{i=0}^{\ell-1}\boldsymbol{a}_{0}^{i},\cdots,\Pi_{i=0}^{\ell-1}\boldsymbol{a}_{k-1}^{i})$. 3. Distribute the shares of these these products $\llbracket\boldsymbol{a}^{\prime i}\rrbracket_{d},i\in[\ell]$ to the parties.
    Functionality $\mathcal{F}_{PreOR}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}^{0}\rrbracket_{d},\cdots,$ $\llbracket\boldsymbol{a}^{\ell-1}\rrbracket_{d},$ where $\boldsymbol{a}^{i}$ is a bit vector of length $k$, $\boldsymbol{a}^{i}=(\boldsymbol{a}_{0}^{i},\cdots,\boldsymbol{a}_{k-1}^{i}),\boldsymbol{a}^{i}_{j}\in\{0,1\}.$ Output: Compute the following 1. Reconstruct the bitwise decomposition of $k$ values $\boldsymbol{a}_{j},j\in[k]$ such that $\boldsymbol{a}_{j}=\sum_{i=0}^{\ell-1}2^{i}\boldsymbol{a}_{j}^{i}$. 2. Compute all prefix-ORs $\boldsymbol{a}^{\prime 0}=(\boldsymbol{a}_{0}^{0},\cdots,\boldsymbol{a}_{k-1}^{0}),$ $\boldsymbol{a}^{\prime 1}=(\boldsymbol{a}_{0}^{0}\vee\boldsymbol{a}_{0}^{1},\cdots,\boldsymbol{a}_{k-1}^{0}\vee\boldsymbol{a}_{k-1}^{1})$, $\boldsymbol{a}^{\prime\ell-1}=(\vee_{i=0}^{\ell-1}\boldsymbol{a}_{0}^{i},\cdots,\vee_{i=0}^{\ell-1}\boldsymbol{a}_{k-1}^{i})$. 3. Distribute the shares of these prefix-ORs $\llbracket\boldsymbol{a}^{\prime i}\rrbracket_{d},i\in[\ell]$ to the parties.
[…]
    Functionality $\mathcal{F}_{Bitwise-LT}$ Input: The functionality receives inputs $k$ values $\boldsymbol{a}_{j},j\in[k]$ and $\ell$ shares $\llbracket\boldsymbol{b}^{i}\rrbracket_{d},i\in[\ell]$ where $\boldsymbol{b}^{i}=\{0,1\}^{k},\boldsymbol{b}_{j}=\sum_{i=0}^{\ell}2^{i}\boldsymbol{b}_{j}^{i}.$ Output: Compute the following 1. Reconstruct $\boldsymbol{b}^{i}$ to compute $\boldsymbol{b}_{j}=\sum_{i=0}^{\ell-1}2^{i}\boldsymbol{b}_{j}^{i},j\in[k].$ 2. Compute $\boldsymbol{c}_{j}=(\boldsymbol{a}_{j}<\boldsymbol{b}_{j}),$ where $c\in\{0,1\}.$ 3. Distribute the shares of $\llbracket\boldsymbol{c}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{DReLU}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}\rrbracket_{d}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}=(\boldsymbol{a}_{0},\cdots,\boldsymbol{a}_{k-1})$ to compute $\boldsymbol{b}_{i}=\text{DReLU}(\boldsymbol{a}_{i}),i\in[k]$. 2. Distribute the shares of $\llbracket\boldsymbol{b}\rrbracket_{d}$ to the parties.
    -B The protocol $\Pi_{Random}$$\Pi_{RandomBits}$$\Pi_{DegreeTrans}$$\Pi_{Xor}$
    -B1 $\Pi_{Random}$
    Batch generation of random shares using Vandermonde matrices. The protocol implementation in Algorithm 13. It is a natural extension of protocols for generating random shares based on Shamir secret sharing. Note that the degree of the polynomial chosen in Algorithm 13 is $d$. We can select any degree $d_{1}$ as long as it satisfies $d\leq d_{1}\leq n-1$. The amortized communication complexity per generated share is 1 round with $\frac{n-1}{k(n-t)}$ field elements.
    Input: None.
    Output: $\llbracket\boldsymbol{r}_{1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{r}_{n-t}\rrbracket_{d}$.
    1. 1.
      Each party $P_{i}$ chooses a random vector $\boldsymbol{u}^{i}\in\mathbb{F}_{p}^{k}$ and distributs $\llbracket\boldsymbol{u}^{i}\rrbracket_{d}$ to other parties.
    2. 2.
      Holding $n$ shares $(\llbracket\boldsymbol{u}^{1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{u}^{n}\rrbracket_{d}),$ each party $P_{i}$ computes $(\llbracket\boldsymbol{r}_{1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{r}_{n-t}\rrbracket_{d})=Van(n,n-t)^{T}\cdot(\llbracket\boldsymbol{u}^{1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{u}^{n}\rrbracket_{d})$.
    return $\llbracket\boldsymbol{r}_{1}\rrbracket_{d},\cdots,\llbracket\boldsymbol{r}_{n-t}\rrbracket_{d}$.
    Algorithm 13 Random Shares $\Pi_{Random}$
    -B2 $\Pi_{DegreeTrans}$
    [3] designed a sharing transformation protocol to securely perform the degree transformation. Specifically, given a PSS $\llbracket\boldsymbol{x}\rrbracket_{n-1}$, it outputs $\llbracket\boldsymbol{x}\rrbracket_{n-k}.$ We extend this protocol simply to support arbitrary degree conversions for Shamir secret sharing and PSS in Algorithm 14. Note that Step 1 can be performed during the offline phase. The amortized communication complexity of the protocol is 2 round with $\frac{2(n-1)}{k}(\frac{1}{n-t}+\frac{1}{n})$ field elements (resp. $2(n-1)(\frac{1}{n-t}+\frac{1}{n})$ field elements for Shamir secret sharing) per party.
    Input: $\llbracket\boldsymbol{x}\rrbracket_{d_{1}}$ (reps. $x|_{s}]_{d_{2}}$), $d\leq d_{1}\leq n-1,d-k+1\leq d_{2}\leq n-1$.
    Output: $\llbracket\boldsymbol{x}\rrbracket_{d_{3}}$ (reps. $x|_{s}]_{d_{4}}$), $d\leq d_{3}\leq n-1,d-k+1\leq d_{4}\leq n-1,d_{1}\neq d_{3},d_{2}\neq d_{4}$.
    1. 1.
[…]
    2. 2.
      All parties compute $\llbracket\boldsymbol{x^{\prime}}\rrbracket_{d_{1}}=\llbracket\boldsymbol{x}\rrbracket_{d_{1}}+\llbracket\boldsymbol{r}\rrbracket_{d_{1}}$ (resp. $x^{\prime}|_{s}]_{d_{2}}=[x|_{s}]_{d_{2}}+[r|_{s}]_{d_{2}}$) and send their shares to $P_{1}$.
    3. 3.
      $P_{1}$ reconstructs the secret $\boldsymbol{x}^{\prime}$ (resp. $x^{\prime}$) and distributes $\llbracket\boldsymbol{x}^{\prime}\rrbracket_{d_{3}}$ (resp. $x^{\prime}|_{s}]_{d_{4}}$) to all parties.
    4. 4.
      All parties locally compute $\llbracket\boldsymbol{x}\rrbracket_{d_{3}}=\llbracket\boldsymbol{x}^{\prime}\rrbracket_{d_{3}}-\llbracket\boldsymbol{r}\rrbracket_{d_{3}}$ (resp. $x|_{s}]_{d_{4}}=[x^{\prime}|_{s}]_{d_{4}}+[r|_{s}]_{d_{4}}$).
    return $\llbracket\boldsymbol{x}\rrbracket_{d_{3}}$ (reps. $x|_{s}]_{d_{4}}$).
    Algorithm 14 Degree Trans $\Pi_{DegreeTrans}$
    -B3 $\Pi_{RandomBits}$
    [35] proposed a protocol to securely generate the shares of random bits. The parties generate a bits $r\in\{0,1\}$ by computing $S(a)=\frac{\frac{a}{\sqrt{a^{2}}}+1}{2}$, where $a\in\mathbb{F}_{p}^{*}$ is a uniformly random non-zero element. $\sqrt{a^{2}}$ is formulated to equal the unique element in [1,(p1)/2][1,(p-1)/2] and $a/\sqrt{a^{2}}$ is $1$ or 1-1. We extend this protocol using PSS in Algorithm 15. The amortized communication complexity per generated random bits is 4 rounds with $\frac{n-1}{k}(\frac{5}{n-t}+\frac{4}{n}+k+1)$ field elements.
    Input: None.
    Output: $\llbracket\boldsymbol{r}\rrbracket_{d},\boldsymbol{r}=\{0,1\}^{k}$.
    1. 1.
      Get $\llbracket\boldsymbol{a}\rrbracket_{d}$ by functionality $\mathcal{F}_{Random}$.
    2. 2.
      Compute $\llbracket\boldsymbol{a}^{\prime}\rrbracket_{d}\leftarrow\mathcal{F}_{Pmult-DN}(\llbracket\boldsymbol{a}\rrbracket_{d},\llbracket\boldsymbol{a}\rrbracket_{d})$.
    3. 3.
      All parties open $\boldsymbol{a}^{\prime}$. If there exists a zero in $\boldsymbol{a}^{\prime}$, then abort. Otherwise, proceed as below.
    4. 4.
      Compute $\boldsymbol{b}_{i}=\sqrt{\boldsymbol{a}^{\prime}_{i}},i\in[k],\llbracket\boldsymbol{b}\rrbracket_{k-1}$ and $\llbracket\boldsymbol{c}\rrbracket_{d+k-1}=\llbracket\boldsymbol{b}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{a}\rrbracket_{d}$ locally.
    5. 5.
      Get $\llbracket\boldsymbol{c}\rrbracket_{d}\leftarrow\mathcal{F}_{DegreeTrans}(\llbracket\boldsymbol{c}\rrbracket_{d+k-1})$.
    6. 6.
      Compute $\llbracket\boldsymbol{r}\rrbracket_{d}=\frac{\llbracket\boldsymbol{c}\rrbracket_{d}+1}{2}.$
    return $\llbracket\boldsymbol{r}\rrbracket_{d}$.
    Algorithm 15 Random Bits $\Pi_{RandomBits}$
    -B4 $\Pi_{Xor}$
    Given two inputs $\llbracket\boldsymbol{a}\rrbracket_{d_{1}},\llbracket\boldsymbol{b}\rrbracket_{d},k-1\leq d_{1}\leq d$, we get $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\oplus\llbracket\boldsymbol{b}\rrbracket_{d}$ by computing $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\oplus\llbracket\boldsymbol{b}\rrbracket_{d}=\llbracket\boldsymbol{a}\rrbracket_{d_{1}}+\llbracket\boldsymbol{b}\rrbracket_{d}-2\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\llbracket\boldsymbol{b}\rrbracket_{d}$, where $\boldsymbol{a},\boldsymbol{b}$ are bit strings. The protocol implementation in Algorithm 16. The amortized communication complexity per bit xor is the same as protocol $\Pi_{DegreeTrans}$.
    Input: $\llbracket\boldsymbol{a}\rrbracket_{d_{1}},\llbracket\boldsymbol{b}\rrbracket_{d},k-1\leq d_{1}\leq d,\boldsymbol{a},\boldsymbol{b}\in\{0,1\}^{k}$.
    Output: $\llbracket\boldsymbol{a}\oplus\boldsymbol{b}\rrbracket_{d},\boldsymbol{a}\oplus\boldsymbol{b}=(\boldsymbol{a}_{0}\oplus\boldsymbol{b}_{0},\cdots,\boldsymbol{a}_{k-1}\oplus\boldsymbol{b}_{k-1})$.
    1. 1.
      Compute $\llbracket\boldsymbol{c}\rrbracket_{d+d_{1}}=2\cdot\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\cdot\llbracket\boldsymbol{b}\rrbracket_{d}.$
    2. 2.
      Set $\llbracket\boldsymbol{a}\oplus\boldsymbol{b}\rrbracket_{d+d_{1}}=\llbracket\boldsymbol{a}\rrbracket_{d_{1}}+\llbracket\boldsymbol{b}\rrbracket_{d}+\llbracket\boldsymbol{c}\rrbracket_{d+d_{1}}.$
    3. 3.
      Get $\llbracket\boldsymbol{a}\oplus\boldsymbol{b}\rrbracket_{d}\leftarrow\mathcal{F}_{DegreeTrans}(\llbracket\boldsymbol{a}\oplus\boldsymbol{b}\rrbracket_{d+d_{1}})$.
    return $\llbracket\boldsymbol{a}\oplus\boldsymbol{b}\rrbracket_{d}$.
    Algorithm 16 Xor $\Pi_{Xor}$
    -C Proofs
    Theorem 13 (Theorem 4).
    A Shamir secret share $x|_{a}]_{d},$ where $f(a)=x,n=2d+1,a\notin\{1,2,\cdots,n\},f$ is a degree-$d$ polynomial. Each of the $n$ parties holds its own share $f(i),i\in\{1,2,\cdots,n\}$. The secret value $x$ can be locally converted to $x|_{b}]_{2d}$, with the conversion performed as $f^{\prime}(i)=f(i)\Pi_{j=1,j\neq i}^{n}\frac{a-j}{b-j}$.
    Proof.
    $x$ $=\sum_{i=1}^{n}f(i)\Pi_{j=1,j\neq i}^{n}\frac{a-j}{i-j}=\sum_{i=1}^{t+1}f(i)\Pi_{j=1,j\neq i}^{t+1}\frac{a-j}{i-j}$ (1)
    $=\sum_{i=1}^{n}(\Pi_{j=1,j\neq i}^{n}\frac{b-j}{i-j})(f(i)\Pi^{n}_{j=1,j\neq i}\frac{a-j}{b-j})$ (2)
    $=\sum_{i=1}^{n}(f(i)\Pi_{j=1,j\neq i}^{n}\frac{b-j}{i-j})(\Pi_{j=1,j\neq i}^{n}\frac{a-j}{b-j})=x$ (3)
    For security proof, we only describe the simulation for protocol $\Pi_{VecMatMult}^{Trunc}$. The security proof for other protocols are simple compositions of local computations and invocations of ideal functionalities.
    Definition 1 (Semi-Honest Security).
    Let $\Pi$ be a $n$-party protocol running in real-world and $\mathcal{F}:(\{0,1\}^{\lambda_{1}})^{n}\rightarrow(\{0,1\}^{\lambda_{2}})^{n}$ be the ideal randomized functionality. We say $\Pi$ securely computes $\mathcal{F}$ in the presence of a subset $\mathcal{C}\subset\{P_{1},P_{2},\cdots,P_{n}\}$ of the parties can be corrupted by semi-honest adversaries if for every corrupted $P_{i},i\in\mathcal{C}$ and every input $\boldsymbol{x}\in(\{0,1\}^{\lambda_{1}})^{n},$ there exists a simulator $\mathcal{S}$ such that:
    $\{view_{i,\Pi}(\boldsymbol{x}),output_{\Pi}(\boldsymbol{x})\}\overset{c}{\approx}\{\mathcal{S}(\boldsymbol{x}_{i},\mathcal{F}_{i}(\boldsymbol{x})),\mathcal{F}(\boldsymbol{x})\}$,
    where $view_{i,\Pi}(\boldsymbol{x})$ is the view of $P_{i}$ in the execution of $\Pi$ on $\boldsymbol{x}$, $output_{\Pi}(\boldsymbol{x})$ is the output of all parties and $\mathcal{F}_{i}(\boldsymbol{x})$ denotes the $i$-th output of $\mathcal{F}(\boldsymbol{x}).$
    Theorem 14 (Theorem 3).
    $\Pi_{VecMatMult}^{Trunc}$ securely realizes $\mathcal{F}_{VecMatMult}^{Trunc}$ in the $\mathcal{F}_{TruncTriple}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    We will construct a simulator $\mathcal{S}$ to simulate the behaviors of honest parties. Let $\mathcal{C}$ denote the set of corrupted parties and $\mathcal{H}$ denote the set of honest parties. The simulator $\mathcal{S}$ works as follows.
    In the beginning, $\mathcal{S}$ receives the input shares of $\llbracket\boldsymbol{a}\rrbracket_{d}$ and $\llbracket A\rrbracket_{d}$ held by corrupted parties. When invoking $\mathcal{F}_{TruncTriple}$, $\mathcal{S}$ invokes the simulator of $\mathcal{F}_{TruncTriple}$ and receives from the adversary the shares of $\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{v},\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}\in\mathbb{F}_{p}^{\lceil v/k\rceil}$ held by corrupted parties.
    In Step 1, for each honest party, $\mathcal{S}$ samples a random vector as its shares of $\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{2d}.$ For each corrupted party, S computes its share of $\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{2d}.$
    In Step 2, depending on whether $P_{1}$ is a corrupted party, there are two cases:
    • If $P_{1}$ is an honest party, $\mathcal{S}$ uses these shares to reconstruct the secret $\boldsymbol{z}=\boldsymbol{a}\cdot A+\boldsymbol{r}$, and sends $\boldsymbol{z}$ back to corrupted parties.
    • If $P_{1}$ is a corrupted party, $\mathcal{S}$ sends the shares of $\boldsymbol{z}=\boldsymbol{a}\cdot A+\boldsymbol{r}$ of honest parties to $P_{1}$. $P_{1}$ also receives the shares from corrupted parties. Then $P_{1}$ can reconstruct $\boldsymbol{z}$ and send $\boldsymbol{z}$ back to corrupted parties.
    In Step 3, $\mathcal{S}$ computes the shares of $\boldsymbol{c}$ held by corrupted parties. Note that $\mathcal{S}$ has computed the shares of $\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$ held by corrupted parties when simulating $\mathcal{F}_{TruncTriple}$.
    In the real view, $\boldsymbol{z}$ is uniformly random since $\boldsymbol{r}$ is uniformly random. Hence, the view generated by the simulator $\mathcal{S}$ is indistinguishable from the real view underlying the security of $\mathcal{F}_{TruncTriple}$. ∎